### v8

# What Do We Get from Two-Way Fixed Effects Regressions? Implications from Numerical Equivalence

Abstract
In any multiperiod panel, a two-way fixed effects (TWFE) regression is numerically equivalent to a first-difference (FD) regression that pools all possible difference lengths. Building on this observation, this paper develops numerical and causal interpretations of the TWFE coefficient. At the sample level, the TWFE coefficient is a weighted average of FD coefficients with varying horizons, which helps clarify the relative contributions of short-run and long-run changes to the overall estimate. At the population level, causal interpretation of the TWFE coefficient relies on a common trends assumption for any time horizons, conditional on changes, not levels, of time-varying covariates. I develop diagnostic procedures to assess plausibility of this assumption across different time horizons.
1 Introduction
Linear regression methods are widely used in empirical economics, valued for their simplicity in translating data into estimates. Yet their ability to deliver clear causal insights is often debated, particularly when treatment effects are heterogeneous (Angrist and Pischke, 2010; Heckman and Urzua, 2010). Two-way fixed effects (TWFE) regressions, increasingly popular in panel data settings, illustrate this tension. They build on linear regression’s intuitive framework to account for unit and time heterogeneity, drawing on the difference-in-differences (DID) framework from the canonical two-period binary treatment setting. However, their use in multiperiod panels with nonbinary treatments, a more complex setting, raises questions about how they map data into coefficients and what causal parameters they capture.
This paper examines the TWFE estimator from both numerical and causal perspectives, without assuming that the motivating linear equation fully reflects the true causal relationships. In panel data with units $i=1,\ldots,N$ and periods $t=1,\ldots,T$, TWFE regressions are based on the equation
$Y_{it}=\alpha_{i}+X_{it}^{\prime}\beta+C_{i}^{\prime}\gamma_{t}+\mu_{t}+\varepsilon_{it},$ (1)
where $Y_{it}$ is a scalar outcome, $\alpha_{i}$ is a unit-specific effect, $X_{it}$ is a vector of explanatory variables, $C_{i}$ is a vector of time-invariant covariates, $\mu_{t}$ is a period-specific effect, and $\varepsilon_{it}$ is a residual. My analysis builds on a simple yet powerful insight: TWFE regressions can be understood through their equivalence to pooled first difference (FD) regressions. Specifically, consider time-series differences of the TWFE equation:
$\Delta_{k}Y_{it}=\Delta_{k}X_{it}^{\prime}\beta+C_{i}^{\prime}\Delta_{k}\gamma_{t}+\Delta_{k}\mu_{t}+\Delta_{k}\varepsilon_{it}\thinspace\thinspace\thinspace\text{for }k=1,\ldots,T-1,$ (2)
where $\Delta_{k}a_{t}\equiv a_{t+k}-a_{t}$ denotes a $k$-period difference of any time series $\{a_{t}\}_{t=1}^{T}$. A least-squares estimate of the TWFE equation (1) and a pooled least-squares estimate of the FD equation (2) across all $k=1,\ldots,T-1$ yield algebraically identical estimates of the coefficient $\beta$. [^2] In standard econometric terminology, "first" in FD refers to the order of differencing (differencing once, as opposed to twice in second differences) rather than the time gap. Thus, this paper uses FD to denote $\Delta_{k}$ operations regardless of the value of $k$. This equivalence, which generalizes the well-known TWFE–FD identity in two-period panels $(T=2$) to multiperiod panel settings ($T\geq 2$), reveals that TWFE regressions capture associations among changes in variables, analogous to how cross-sectional regressions capture associations among levels of variables.
The numerical equivalence between TWFE and pooled FD regressions is not a surprising or theoretically novel result. It is a direct consequence of a basic identity: the sample covariance of two variables equals (up to scale) the average product of all pairwise differences—a well-known fact in the theory of U-statistics. Han and Lee (2017, 2022) apply this identity to the within estimator and express it in terms of pairwise differences. They use this formulation to study bias correction, under the assumption of a correctly specified linear causal model. My contribution is not to rediscover this identity but to reinterpret its implications: I use it to understand the properties of the TWFE estimator without assuming that the linear model is causally valid. While Goodman-Bacon (2021) and Li and Strezhnev (2024) derive related decomposition results in the case of binary regressors, they do so without explicitly connecting their approach to this fundamental statistical identity. Their findings can in fact be seen as special cases of the same underlying numerical structure, though focused specifically on binary treatment settings. By contrast, I make this identity explicit and use it to characterize the TWFE estimator more generally—across binary, discrete, and continuous regressors, and regardless of treatment timing or the inclusion of covariates.
Building on this equivalence, I first establish that the TWFE estimator can be expressed as a weighted average of the FD estimators across all possible between-period gaps. Let $\widehat{\beta}_{\text{FE}}$ be the coefficient on $X_{it}$ from a least-squares estimate of equation (1), and $\widehat{\beta}_{\text{FD},k}$ the coefficient on $\Delta_{k}X_{it}$ from separate least-squares estimates of equation (2) for each $k=1,\ldots,T-1$. It is natural to expect a connection among them, since $\widehat{\beta}_{\text{FE}}$ also arises from a joint least-squares estimate of equation (2) using all $k=1,\ldots,T-1$. In fact, when $X_{it}$ is univariate, $\widehat{\beta}_{\text{FE}}$ is a positively-weighted average of $\left\{\widehat{\beta}_{\text{FD},k}\right\}_{k=1}^{T-1}$:
$\widehat{\beta}_{\text{FE}}=\sum_{k=1}^{T-1}\widehat{w}_{k}\widehat{\beta}_{\text{FD},k},$ (3)
where $\widehat{w}_{k}\geq 0$ and $\sum_{k=1}^{T-1}\widehat{w}_{k}=1$. This decomposition provides a diagnostic tool for understanding how a TWFE regression aggregates short- and long-run associations between changes in $Y_{it}$ and changes in $X_{it}$ into a single coefficient. If $X_{it}$ consists of a scalar treatment variable and time-varying covariates, the weighted-average interpretation remains valid for the coefficient on the treatment, although it relies on an adjustment for differential covariate impacts across periods.
I then develop causal interpretation of TWFE coefficients within a potential outcome framework when $X_{it}$ consists of a scalar treatment ($D_{it}$) and time-varying covariates ($W_{it}$). A key condition for causal interpretation is the conditional common trends assumption:
$E\left[\Delta_{k}Y_{it}(d)|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d)|\Delta_{k}W_{it},C_{i}\right]\thinspace\thinspace\thinspace\text{for }k=1,\ldots,T-1,$
imposing the conditional mean independence of the potential outcome change $\Delta_{k}Y_{it}(d)$, evaluated at treatment level $d$, from the treatment change $\Delta_{k}D_{it}$. Crucially, this condition builds on $k$-period changes as a natural consequence of the revelation that TWFE relies on $k$-period changes. This assumption highlights potential challenges in supporting causal interpretation of the TWFE coefficient, including the difficulty of justifying common trends over increasingly long horizons, and the reliance on conditioning on covariate changes rather than predetermined levels, both contrasting with the canonical two-period DID framework.
The core issue with TWFE lies in its aggregation across all possible difference lengths. I show that individual $k$-period FD regressions can achieve similar causal interpretation under the same identifying conditions, but these conditions need only hold for the specific $k$ rather than across all possible $k$. Traditional panel data econometrics justifies TWFE’s aggregation on efficiency grounds, which rely on three joint conditions: common trends assumptions hold for all difference lengths $k$, treatment effects are static rather than dynamic, and treatment effects are homogeneous. When at least one of these conditions fails, the TWFE aggregation loses its advantage. Researchers may then be better served by $k$-period FD regressions or heterogeneity-robust estimators developed in the recent literature. Since these alternatives vary in their common trends duration requirements, I develop diagnostic procedures that signal violations of the common trends assumption for increasing $k$ using testable implications. These diagnostics help researchers identify the range of difference lengths $k$ over which their identifying assumptions remain plausible and select appropriate methods for their setting.
This paper contributes to the recent literature on TWFE regressions, which has primarily focused on settings with binary treatments. A growing number of studies investigate numerical and causal properties of TWFE estimators with univariate or multivariate binary regressors, diagnose their issues, and propose alternative estimators. The literature covers a range of settings, including binary and staggered treatment cases (e.g., Athey and Imbens, 2022; Callaway and Sant’Anna, 2021; Goodman-Bacon, 2021; Wooldridge, 2021), binary and nonstaggered cases (e.g., de Chaisemartin and D’Haultfœuille, 2020), event-study settings (e.g., Borusyak et al., 2024; Lin and Zhang, 2022; Schmidheiny and Siegloch, 2023; Sun and Abraham, 2020), cases with multiple binary treatment variables (e.g., de Chaisemartin and D’Haultfœuille, 2023), and continuous and staggered treatment settings (e.g., Callaway et al., 2024). While these existing studies offer valuable insights into specific settings, this paper provides general results for TWFE regressions that hold across a range of specifications, including binary, discrete, or continuous regressors, with or without covariates, and under varying treatment dynamics.
The rest of this paper is organized as follows. Section 2 investigates the numerical properties of TWFE regressions, and Section 3 offers causal interpretation of TWFE coefficients. Section 4 discusses the practical implications of these findings and develops diagnostic procedures for assessing common trends violations across different time horizons. Section 5 illustrates the practical relevance of these insights by examining the TWFE estimates of the minimum wage effect on employment outcomes. Section 6 concludes. All proofs are in Appendix A.
2 Numerical Properties
This section develops a numerical interpretation of the TWFE estimator. I focus on a balanced panel for simplicity; Appendix C discusses how the results in the main paper extend to or differ in an unbalanced panel. For any time series $\{a_{t}\}_{t=1}^{T}$, I use $\Delta_{k}a_{t}\equiv a_{t+k}-a_{t}$ to denote a $k$-period difference and $\overline{a}\equiv\frac{1}{T}\sum_{t=1}^{T}a_{t}$ to denote the time-series average.
2.1 Equivalence of Least-Square Objectives
In a two-period panel, it is well known that TWFE and FD regressions yield the same coefficient estimates. This equivalence extends to multiperiod panels as well. The following theorem shows that the TWFE estimator can be obtained from a least-squares problem that pools across all $k$-period first differences. This result holds for both univariate and multivariate regressors, whether they are binary, discrete, or continuous.
Theorem 1.
Let $\widehat{\beta}_{\text{FE}}$ be the coefficient on $X_{it}$ from a least-squares problem:
$\underset{\beta,\{\alpha_{i}\}_{i=1}^{N},\{\gamma_{t},\mu_{t}\}_{t=1}^{T}}{\min}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}-\alpha_{i}-X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{t}-\mu_{t}\right)^{2}.$ (4)
Then $\beta=\widehat{\beta}_{\text{FE}}$ is also solves:
$\underset{\beta,\{\gamma_{t},\mu_{t}\}_{t=1}^{T}}{\min}$ $\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\Delta_{k}\gamma_{t}-\Delta_{k}\mu_{t}\right)^{2}.$ (5)
The equivalence follows from a fundamental statistical identity: the sample covariance between $\{a_{t}\}_{t=1}^{T}$ and $\{b_{t}\}_{t=1}^{T}$ equals the average of ($a_{s}-a_{t})(b_{s}-b_{t})$ over all pairs with $s>t$, up to normalization. This identity—closely related to U-statistics—is what underlies the equivalence between the TWFE and pooled FD objectives. Han and Lee (2017, 2022) leverage this identity in a setting without period-specific parameters $\{\gamma_{t},\mu_{t}\}_{t=1}^{T}$ in (4), which accommodates TWFE regressions when $X_{it}$ includes time dummies and their interactions with $C_{i}$. They show that the within estimator, given by a regression of $Y_{it}-\overline{Y}_{i}$ on $X_{it}-\overline{X}_{i}$, can alternatively be recovered via a regression of $\Delta_{k}Y_{it}$ on $\Delta_{k}X_{it}$, and use this result to study bias correction, assuming a correctly specified linear model. My approach, by contrast, applies this numerical property to reinterpret TWFE regressions without assuming a linear causal model.
The pooled FD objective reveals that TWFE estimates reflect changes in the variables of interest across all possible time differences (from one period to T1T-1 periods), rather than their levels. This shift in perspective opens a door to interpreting TWFE estimators in much broader settings than previously emphasized. For binary treatment settings, Goodman-Bacon (2021) and Li and Strezhnev (2024) show that TWFE estimators can be expressed as averages of two-unit, two-period DID comparisons. While their analyses, tailored to these settings, do not explicitly invoke the statistical identity used here, my approach reveals that their results stem from the same fundamental numerical structure. Appendix D elaborates on this connection.
2.2 Weighted-Average Relationship
A least-squares estimate of equation (2) using only $k$-period differences is given by:
$\underset{\beta,\{\gamma_{t},\mu_{t}\}{}_{t=1}^{T}}{\min}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\Delta_{k}\gamma_{t}-\Delta_{k}\mu_{t}\right)^{2},$ (6)
[…]
$\widetilde{Y}_{it}\equiv Y_{it}-\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)^{\prime}\left(\sum_{i=1}^{N}\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)^{\prime}\right)^{-1}\sum_{i=1}^{N}\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)Y_{it}$ (7)
to be a residual from a regression of $Y_{it}$ on $C_{i}$ independently performed for each $t=1,\ldots,T$. Then, applying the Frisch–Waugh–Lovell (FWL) theorem to the two regressions (5) and (6) results in a weighted-average interpretation of the TWFE coefficient.
Theorem 2.
The TWFE coefficient that solves (5) is given by:
$\widehat{\beta}_{\text{FE}}$ $=\left(\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right)^{-1}\left(\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{Y}_{it}\right),$
provided that $\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}$ is positive definite. The k-period FD coefficient $\widehat{\beta}_{\text{FD},k}$ that solves (6) is:
$\widehat{\beta}_{\text{FD},k}=\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right)^{-1}\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{Y}_{it}^{\prime}\right),$
provided that $\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}$ is positive definite. If $\widehat{\beta}_{\text{FE}}$ and $\left\{\widehat{\beta}_{\text{FD},k}\right\}_{k=1}^{T-1}$ are all uniquely defined,
$\widehat{\beta}_{\text{FE}}=\sum_{k=1}^{T-1}\widehat{\varOmega}_{k}\widehat{\beta}_{\text{FD},k},$ (8)
where the weight matrices for $k=1,\ldots,T-1$ are:
$\widehat{\varOmega}_{k}$ $=\left(\sum_{\ell=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-\ell}\Delta_{\ell}\widetilde{X}_{it}\Delta_{\ell}\widetilde{X}_{it}^{\prime}\right)^{-1}\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right),$
which are all positive definite and sum to the identity matrix. If some $\widehat{\beta}_{\text{FD},k}$ are not uniquely defined due to $\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}$ not being full-rank, then there are multiple $\widehat{\beta}_{\text{FD},k}$ that minimizes (6), and for any of them the identity (8) still holds.
When $X_{it}$ is univariate, Theorem 2 implies that $\widehat{\beta}_{\text{FE}}$ is a convex combination of $\widehat{\beta}_{\text{FD},k}$, as presented in (3). When $X_{it}$ is multivariate, Theorem 2 provides a matrix-weighted average interpretation, which may not be very intuitive for practical applications that often focus on the effect of a single treatment variable. Thus, I consider a common scenario in which $X_{it}$ consists of a scalar treatment variable $D_{it}$ that produces a causal effect of interest and a vector of time-varying covariates $W_{it}$, and focus on the interpretation of the coefficient on $D_{it}$. Using $D_{it}$ as the first element of $X_{it}$, I write:
$X_{it}=\left[\begin{array}[]{c}D_{it}\\ W_{it}\end{array}\right],\thinspace\thinspace\thinspace\widehat{\beta}_{\text{FE}}=\left[\begin{array}[]{c}\widehat{\beta}_{\text{FE}}^{D}\\ \widehat{\beta}_{\text{FE}}^{W}\end{array}\right],\thinspace\thinspace\thinspace\text{and}\,\thinspace\thinspace\widehat{\beta}_{\text{FD},k}=\left[\begin{array}[]{c}\widehat{\beta}_{\text{FD},k}^{D}\\ \widehat{\beta}_{\text{FD},k}^{W}\end{array}\right].$
Theorem 2 suggests that $(\widehat{\beta}_{\text{FE}}^{D},\widehat{\beta}_{\text{FE}}^{W})$ is a matrix-weighted average of $(\widehat{\beta}_{\text{FD},k}^{D},\widehat{\beta}_{\text{FD},k}^{W})$. This implies that $\widehat{\beta}_{\text{FE}}^{D}$ depends on both $\widehat{\beta}_{\text{FD},k}^{D}$ and $\widehat{\beta}_{\text{FD},k}^{W}$ unless the weight matrix is diagonal. The following theorem formally shows that $\widehat{\beta}_{\text{FE}}^{D}$ can be expressed as a weighted average of $\widehat{\beta}_{\text{FD},k}^{D}$ after an adjustment for the influence of time-varying covariates.
Theorem 3.
The TWFE coefficient on $D_{it}$ is given by:
$\widehat{\beta}_{\text{FE}}^{D}=\sum_{k=1}^{T-1}\widehat{w}_{k}\widehat{\beta}_{\text{FD},k}^{D}+\frac{\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)^{\prime}\left(\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right)\left(\widehat{\beta}_{\text{FD},k}^{W}-\widehat{\beta}_{\text{FE}}^{W}\right)}{\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{D}_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)},$ (9)
where $\widehat{\delta}_{\text{FE}}^{W}$ and $\widehat{\delta}_{\text{FD},k}^{W}$ are the coefficients from the TWFE and $k$-period FD regressions of $D_{it}$ on $W_{it}$, and the weights are defined as:
$\widehat{w}_{k}\equiv\frac{\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{D}_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)}{\sum_{\ell=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-\ell}\Delta_{\ell}\widetilde{D}_{it}\left(\Delta_{\ell}\widetilde{D}_{it}-\Delta_{\ell}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)}.$ (10)
This theorem shows that $\widehat{\beta}_{\text{FE}}^{D}$ can be decomposed into two components. The first is a weighted average of the FD coefficients. The weights, which reflect variation in treatment changes not explained by covariates or time effects, sum to one but may be negative in exceptional cases. [^4] The denominator in (10) equals the sum of $(\Delta_{\ell}\widetilde{D}_{it}-\Delta_{\ell}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W})^{2}$, while the numerator equals the sum of $(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FD},k}^{W})(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W})$ and is not identical to the sum of $(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W})^{2}$ unless $\widehat{\delta}_{\text{FE}}^{W}=\widehat{\delta}_{\text{FD},k}^{W}$ or $\widehat{\delta}_{\text{FE}}^{W}=0$. As such, some of the weights do not remain positive in highly unusual cases where unexplained treatment changes given by the TWFE and $k$-period FD regressions are negatively correlated. The second is an adjustment term, which arises from discrepancies between the TWFE and FD regressions in how they account for covariate effects. While a TWFE regression applies the same coefficients $(\widehat{\beta}_{\text{FE}}^{W}$ and $\widehat{\delta}_{\text{FE}}^{W}$) to all $k$-period covariate changes, FD regressions allow the coefficients on covariate changes ($\widehat{\beta}_{\text{FD},k}^{W}$ and $\widehat{\delta}_{\text{FD},k}^{W}$) to vary across $k=1,\ldots,T-1$, leading to a deviation from a clean weighted-average interpretation. While the adjustment term complicates this interpretation, it has implications for causal interpretation of the population TWFE coefficient, as discussed in Section 3.
The decomposition in Theorem 3 differs in spirit from existing TWFE decompositions (e.g., Goodman-Bacon, 2021), which typically express regression coefficients as weighted averages of readily interpretable quantities. Instead, this decomposition serves as a diagnostic tool for probing TWFE’s identifying assumptions by revealing how it implicitly aggregates associations across all possible time gaps (k=1,,T1k=1,...,T-1). Systematic variation in $\widehat{\beta}_{\text{FD},k}^{D}$ across different $k$ signals potential violations of key assumptions underlying TWFE’s causal interpretation—whether from failures of common trends for certain time horizons, dynamic treatment effects, or effect heterogeneity. Such patterns encourage researchers to investigate whether TWFE’s implicit aggregation aligns with their economic question and identifying assumptions. Section 3 formalizes this diagnostic perspective by providing causal interpretations of TWFE and FD coefficients under different assumption sets.
3 Causal Interpretation
3.1 TWFE Interpretation
This section develops causal interpretation of the population TWFE coefficient, focusing on the case in which $X_{it}$ consists of a scalar treatment ($D_{it}$) and a vector of time-varying covariates ($W_{it}$). I consider a large–$N$ fixed–$T$ setting, with $\left(Y_{it},D_{it},W_{it}\right)_{t=1}^{T}$ and $C_{i}$ being independent and identically distributed across $i=1,\ldots,N$. The cross-sectional mean is represented by $E\left[Y_{it}\right]$, while the time-series mean is denoted by $\overline{Y}_{i}$. In addition, I define
$\widetilde{Y}_{it}\equiv Y_{it}-\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)^{\prime}E\left[\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)^{\prime}\right]^{-1}E\left[\left(\begin{array}[]{c}1\\ C_{i}\end{array}\right)Y_{it}\right]$
to be a residual from the population projection of $Y_{it}$ on $(1,C_{i})$. This parallels the sample definition in (7); for simplicity of notation I reuse the same symbol.
The equivalence result in Section 2 reduces a TWFE regression to a regression of $\Delta_{k}\widetilde{Y}_{it}$ on $\Delta_{k}\widetilde{D}_{it}$ controlling for $\Delta_{k}\widetilde{W}_{it}$. The population TWFE coefficient on $D_{it}$ is given by
$\beta_{\text{FE}}^{D}=\frac{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]}{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]},$ (11)
where
$\delta_{\text{FE}}^{W}=\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{D}_{it}\right]\right)$ (12)
represents the population version of $\widehat{\delta}_{\text{FE}}^{W}$. With incidental unit-specific effect parameters being eliminated, this structure becomes analogous to a standard cross-sectional regression, making the problem tractable with standard causal inference tools. Consequently, causal interpretation of $\beta_{\text{FE}}^{D}$ can be developed within a potential outcome framework using the following assumptions.
Assumption PO.
(Potential Outcome, Static) For each $t=1,\ldots,T$, $\{Y_{it}(d):d\in(\underline{d},\overline{d})\}$ is a stochastic process that defines a potential outcome associated with each possible treatment level $d\in(\underline{d},\overline{d})$, where $\infty\leq\underline{d}<\overline{d}\leq\infty$. The observed outcome is given by $Y_{it}=Y_{it}(D_{it})$.
Assumption PO rules out dynamic treatment effects by restricting a potential outcome to be a function only of the current treatment status. Dynamic effects are typically allowed in a staggered adoption design setting with a binary treatment (e.g., Callaway and Sant’Anna, 2021; Goodman-Bacon, 2021). In a more general setting considered here, Assumption PO is necessary to avoid making a potential outcome an extremely high-dimensional object. Appendix B.1 relaxes this assumption to allow for dynamic effects, though under a homogeneous treatment effect restriction.
Assumption CT.
(Conditional Common Trends) There exists $d^{0}\in(\underline{d},\overline{d})$ such that
$E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]$
for any $(k,t)$ with $1\leq t<t+k\leq T$.
Assumption CT requires that trends in potential outcomes $\Delta_{k}Y_{it}(d^{0})$ at a baseline treatment level $d^{0}$ be mean independent from treatment changes $\Delta_{k}D_{it}$, conditional on time-invariant covariates $C_{i}$ and the concurrent changes $\Delta_{k}W_{it}$ in time-varying covariates. This assumption presumes the existence of a natural baseline treatment level with economic meaning, such as "no treatment" ($d^{0}=0$) in a binary treatment setting. The baseline level $d^{0}$ serves as the counterfactual reference point for causal interpretation; the TWFE coefficient measures treatment effects relative to this baseline. [^5] One could select any arbitrary $d^{0}$ and assert that the assumption holds. However, doing so creates two problems: first, the resulting causal interpretation becomes meaningless as effects are measured relative to an economically irrelevant reference point; second, the assumption becomes difficult to defend on economic grounds. The baseline level $d^{0}$ need not be experienced in the data as long as it represents an economically meaningful reference point and the parallel trends assumption at $d^{0}$ can be economically justified. [^6] For instance, a tariff rate of zero in trade policy analysis may provide a natural baseline even if no country implements complete free trade, and common trends in the absence of trade barriers may be plausible on economic grounds. When no natural baseline exists, such as with minimum wage policy where all levels represent active intervention, it can be problematic to assume common trends hold at one arbitrarily chosen level while potentially failing at all others. Appendix B.2 addresses this issue by invoking the common trends assumption for every treatment level.
Assumption V.
(Variation) $\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right]>0$.
Assumption V requires that $\Delta_{k}D_{it}$ have some variation conditional on $\left(\Delta_{k}W_{it},C_{i}\right)$.
Assumption L.
(Linearity) For any $(k,t)$ with $1\leq t<t+k\leq T$, the conditional expectation of $\Delta_{k}D_{it}$ given $\left(\Delta_{k}W_{it},C_{i}\right)$ can be expressed as
$E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]=\Delta_{k}W_{it}^{\prime}\delta_{k,t}^{W}+C_{i}^{\prime}\delta_{k,t}^{C}+\delta_{k,t}^{I}.$
Assumption L requires that the conditional mean $E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]$ be linear in $\left(\Delta_{k}W_{it},C_{i}\right)$. Similar linearity assumptions have been commonly used in interpreting linear regression coefficients in the presence of covariates. [^7] See, for example, Angrist and Krueger (1999), Lochner and Moretti (2015), Słoczyński (2020), Ishimaru (2024), and Goldsmith-Pinkham et al. (2024). This assumption simplifies the analysis by ruling out omitted variable bias that arises from unmodeled nonlinear effects of covariates.
Assumption L is not sufficient for enabling causal interpretation of $\beta_{\text{FE}}^{D}$ without any omitted variable bias. Thus, I make the following assumption to eliminate the bias.
Assumption LH.
(Linearity and Time Homogeneity) For any $(k,t)$ with $1\leq t<t+k\leq T$, the conditional expectation of $\Delta_{k}D_{it}$ given $\left(\Delta_{k}W_{it},C_{i}\right)$ can be expressed as
$E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]=\Delta_{k}W_{it}^{\prime}\delta^{W}+C_{i}^{\prime}\delta_{k,t}^{C}+\delta_{k,t}^{I},$
where the coefficient $\delta^{W}$ does not vary across $(k,t)$.
In addition to the linearity requirement made by Assumption L, Assumption LH requires that the relationship between $\Delta_{k}D_{it}$ and $\Delta_{k}W_{it}$ be homogeneous across periods. Assumption L can be made as nonrestrictive as possible by choosing a flexible specification of covariates (e.g., adding polynomial and interaction terms), although small sample size may limit this in practice. In contrast, Assumption LH imposes a fundamental restriction on the relationship between treatment changes and covariate changes, which cannot be relaxed by increasing flexibility within the TWFE framework.
Assumption LH is closely related to an imperfect weighted-average formula in Theorem 3. The following lemma clarifies the relationship.
Lemma 1.
Under Assumption LH, $\delta_{\text{FE}}^{W}=\delta^{W}=\delta_{\text{FD},k}^{W}$ for all $k=1,\ldots,T-1$, where $\delta_{\text{FE}}^{W}$ is as defined in (12) and $\delta_{\text{FD},k}^{W}$ is a coefficient from a regression of $\Delta_{k}\widetilde{D}_{it}$ on $\Delta_{k}\widetilde{W}_{it}$.
Lemma 1 suggests that Assumption LH is a sufficient condition for $\widehat{\delta}_{\text{FE}}^{W}$ and $\widehat{\delta}_{\text{FD},k}^{W}$ to be equivalent in the population. Therefore, under Assumption LH the adjustment term in Theorem 3 vanishes in the population.
Combining all the assumptions yields the following result.
Theorem 4.
Under Assumptions PO, V, L, and CT,
$\beta_{\text{FE}}^{D}=\sum_{t=1}^{T}E\left[\tau_{it}\omega_{it}\right]+B,$
where
$\tau_{it}\equiv\begin{cases}\frac{Y_{it}(D_{it})-Y_{it}(d^{0})}{D_{it}-d^{0}}&D_{it}\neq d^{0}\\ 0&D_{it}=d^{0}\end{cases}$
is a per-unit effect of a deviation of the treatment $D_{it}$ from the baseline level $d^{0}$. The weights satisfy $\sum_{t=1}^{T}E\left[\omega_{it}\right]=1$ and
$\omega_{it}\propto\left(D_{it}-d^{0}\right)\left\{\left(\widetilde{D}_{it}-\overline{\widetilde{D}}_{i}\right)-\left(\widetilde{W}_{it}-\overline{\widetilde{W}}_{i}\right)^{\prime}\delta_{\text{FE}}^{W}\right\}.$
The bias term $B$ is given by
$B=\frac{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d^{0})\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\delta_{k,t}^{W}-\delta_{\text{FE}}^{W}\right)\right]}{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]},$
which satisfies $B=0$ under Assumption LH.
Theorem 4 interprets $\beta_{\text{FE}}^{D}$ as a weighted average of per-unit treatment effects, with weights summing to one but possibly negative. The weight function $\omega_{it}$ interacts a deviation of $D_{it}$ from the baseline level $d^{0}$ with a TWFE residual of $D_{it}$.
However, this weighted-average interpretation faces multiple challenges. Three issues concern the assumptions needed for the weighted-average interpretation, while a fourth issue arises even when this interpretation holds.
First, the common trends assumption (Assumption CT) must hold across all difference lengths $k$. This requirement extends far beyond the canonical two-period DID setting, and becomes increasingly difficult to justify as the number of time periods grows. Section 4 explores practical implications from this requirement in detail.
Second, the role of time-varying covariates in a TWFE regression differs entirely from a two-period DID setting that allows for covariates. When a TWFE regression includes time-varying covariates, a common trends assumption must be made conditional on changes in covariates, not their initial levels, as illustrated by Assumption CT. By contrast, a common trends assumption is typically made conditional on pretreatment observables in a two-period DID setting (Heckman et al., 1997, 1998; Abadie, 2005). As formalized by Caetano et al. (2022), a common trends assumption that uses covariate changes is less theoretically favored than the one that uses predetermined covariate levels. In particular, if the treatment change $\Delta_{k}D_{it}$ influences the covariate change $\Delta_{k}W_{it}$, then variation in $\Delta_{k}D_{it}$ with $\Delta_{k}W_{it}$ being fixed does not correspond to ceteris paribus variation that defines the causal effect of the treatment. [^8] This type of issue is often called a “bad control” problem, a term popularized by Angrist and Pischke (2009, Section 3.2.3). They describe occupation as a “bad control” when examining the effects of a college degree on earnings. Indeed, occupation is not held constant when we typically define the ceteris paribus effect of a college degree on earnings.
Third, the linearity and time homogeneity assumption (Assumption LH) can be restrictive unless all covariates are discrete and time-invariant. This assumption demands that the relationship between treatment changes and covariate changes be not only linear but also constant across all time periods and difference lengths. Despite being essential for causal interpretation, these requirements are typically assumed implicitly rather than explicitly defended, unlike common trends assumptions.
Fourth, even when the weighted-average interpretation holds, the weights $\omega_{it}$ can be negative, as similarly highlighted in a binary treatment case (e.g., Borusyak et al., 2024; de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021). Whether this poses a problem depends on the correlation between weights and treatment effects. If some weights are negative and weights are strongly correlated with per-unit treatment effects, then $\beta_{\text{FE}}^{D}$ can even lie outside the support of per-unit treatment effects $\tau_{it}$. Nevertheless, negative weights alone are not necessarily problematic if treatment effects are uncorrelated with weights, as noted by de Chaisemartin and D’Haultfœuille (2020) in their Corollary 2. Appendix B.2 shows that the weight function $\omega_{it}$ contains redundant variation that is orthogonal to treatment effects under a more strict common trends assumption, which may alleviate the concern about negative weights in some cases.
Overall, these issues highlight that causal interpretation of TWFE estimators involves more restrictive assumptions than commonly recognized, particularly regarding the length over which common trends can be justified, the role of time-varying covariates, and the need for linearity and time homogeneity conditions, all of which are rarely scrutinized in empirical practice.
Remarks
de Chaisemartin and D’Haultfœuille (2020) provide a weighted-average interpretation of the TWFE coefficient that resembles Theorem 4 in their online appendix. They separately consider cases with nonbinary treatment without covariates and binary treatment with time-varying covariates. However, their proof can be straightforwardly extended to allow for nonbinary treatment with both time-invariant and time-varying covariates. In the current setup, their assumption can be expressed as:
$E\left[\Delta_{k}Y_{it}(d^{0})|D_{i1},\ldots,D_{iT},W_{i1},\ldots,W_{iT},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right],$ (13)
which is a sufficient condition for Assumption CT. Note that (13) conditions on entire treatment and covariate histories $(D_{i1},\ldots,D_{iT},W_{i1},\ldots,W_{iT})$, whereas Assumption CT conditions only on concurrent changes $(\Delta_{k}D_{it},\Delta_{k}W_{it})$. They also assume linearity and time homogeneity of $E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]$. This assumption corresponds to Assumption LH, which imposes the same condition on $E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]$. [^9] Linearity and time homogeneity of either $\Delta_{k}Y_{it}(d^{0})$ or $\Delta_{k}D_{it}$ is sufficient to eliminate the bias term $B$ in Theorem 4. Given its connection to the weighted-average numerical interpretation in Theorem 3 (demonstrated by Lemma 1) and its testability with observed data, my analysis focuses on the setting with Assumption LH rather than conditions on potential outcomes.
The contribution of Theorem 4 relative to their result lies in showing that Assumption CT, which relies solely on concurrent changes rather than full histories, is central to causal interpretation of the TWFE coefficient. The importance of covariate changes rather than levels for TWFE interpretation is well-recognized in canonical two-period settings, stemming from the known equivalence between TWFE and FD. Theorem 4 extends this insight to multiperiod settings, demonstrating that concurrent changes remain the fundamental basis for causal interpretation—a non-trivial extension. While Caetano and Callaway (2024) also highlight the importance of covariate changes for TWFE interpretation, their analysis focuses primarily on the two-period case. In extending their approach to multiperiod settings, they—like de Chaisemartin and D’Haultfœuille (2020)—rely on conditioning on full covariate histories, whereas Theorem 4 demonstrates the sufficiency of conditioning only on concurrent changes.
While this emphasis on concurrent changes clarifies the role of covariates, the implications for treatment dynamics warrant closer examination. Exogeneity of an entire treatment series $\left(D_{i1},\ldots,D_{iT}\right)$ as in (13), often called a strict exogeneity condition, is a common assumption in panel data models. This assumption rules out the possibility that a shock to the current outcome influences future treatment status or is correlated with past treatment status. Although constraints that Assumption CT imposes on treatment dynamics appear weaker than strict exogeneity when considered for each $k=1,\ldots,T-1$ individually, similar restrictions arise when considered collectively. Specifically, the common trends assumption for periods $t$ to $t+2k$ inherently makes it unlikely for potential outcome changes $\Delta_{k}Y_{it}(d)$ in the first $k$ periods to influence treatment changes $\Delta_{k}D_{i,t+k}$ in the subsequent $k$ periods. It also makes correlations between prior treatment changes $\Delta_{k}D_{it}$ and subsequent potential outcome changes $\Delta_{k}Y_{i,t+k}(d)$ practically implausible. Without excluding such scenarios, correlations between $\Delta_{2k}Y_{it}(d)$ and $\Delta_{2k}D_{it}$ would arise, even in the absence of correlations between $\Delta_{k}Y_{it}(d)$ and $\Delta_{k}D_{it}$ or between $\Delta_{k}Y_{i,t+k}(d)$ and $\Delta_{k}D_{i,t+k}$. This consideration is essential for understanding the potential limitations of the common trends assumption over longer time horizons, as further discussed in Section 4.
3.2 FD Interpretation
The causal interpretation of the $k$-period FD coefficient follows directly from the TWFE framework, but requires assumptions only for the specific difference length $k$:
Assumption CT–k.
(Conditional Common Trends, $k$-period FD) There exists $d^{0}\in(\underline{d},\overline{d})$ such that the following holds for any $t\in\{1,\ldots,T-k\}$:
$E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right].$
Assumption V–k.
(Variation, $k$-period FD) $\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right]>0$.
Assumption LH–k.
(Linearity and Time Homogeneity, $k$-period FD) For $t\in\{1,\ldots,T-k\}$, the conditional expectation of $\Delta_{k}D_{it}$ given $\left(\Delta_{k}W_{it},C_{i}\right)$ can be expressed as
$E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]=\Delta_{k}W_{it}^{\prime}\delta_{k}^{W}+C_{i}^{\prime}\delta_{k,t}^{C}+\delta_{k,t}^{I},$
where the coefficient $\delta_{k}^{W}$ does not vary across $t$.
These assumptions are identical to Assumptions CT, V, and LH, except that they need only hold for the specific $k$ rather than across all $k=1,\ldots,T-1$. For simplicity, I directly invoke the linearity and time homogeneity assumption rather than building up from linearity alone.
The population $k$-period FD coefficient $\beta_{\text{FD},k}^{D}$ is obtained from a regression of $\Delta_{k}Y_{it}$ on $\Delta_{k}D_{it}$ controlling for $\Delta_{k}W_{it}$ and $C_{i}$ interacted with period indicators. Under the stated assumptions, this coefficient has the following causal interpretation.
Theorem 5.
Under Assumptions PO, CT–k, V–k, and LH–k,
$\beta_{\text{FD},k}^{D}=\sum_{t=1}^{T}E\left[\tau_{it}\omega_{it}^{(k)}\right],$
where $\tau_{it}$ is a per-unit effect defined in Theorem 4. The weights satisfy $\sum_{t=1}^{T}E\left[\omega_{it}^{(k)}\right]=1$ and
$\omega_{it}^{(k)}\propto(D_{it}-d^{0})\sum_{s=1}^{T}\mathbbm{1}_{|t-s|=k}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{k}^{W}\right).$
Theorem 5 suggests that the $k$-period FD coefficient admits a weighted-average interpretation of per-unit treatment effects, with weights that depend on the unpredictable component of $k$-period treatment changes.
However, the FD interpretation inherits several problems from the TWFE framework. The common trends assumption remains difficult to justify for large $k$, though researchers can mitigate this concern by choosing smaller values of $k$. The identification strategy still relies on covariate changes rather than predetermined covariate levels, creating the same theoretical concerns about ceteris paribus variation. Covariate impacts should be homogeneous across time periods $t$, though homogeneity across difference lengths $k$ is no longer required. Finally, the possibility of negative weights persists, complicating the interpretation of the weighted average.
Among these inherited problems, those associated with covariates can be readily addressed within the FD framework. Rather than controlling for covariate changes $\Delta_{k}W_{it}$, FD regressions can be modified to control for the initial covariate levels $W_{it}$ or other predetermined variables. Moreover, the homogeneity restriction on covariate impacts across time periods can be relaxed by allowing period-specific coefficients.
The FD framework thus offers greater flexibility than the TWFE approach in addressing covariate-related concerns while requiring weaker assumptions overall. Given that a TWFE regression is merely a pooled version of FD regressions as shown in Section 2, this raises the question of whether such aggregation is desirable when an individual FD coefficient may admit causal interpretation under more plausible conditions.
4 Practical Implications and Diagnostics
4.1 Discussion: When Should We Use TWFE?
Both TWFE and $k$-period FD coefficients admit weighted-average interpretations of causal effects. However, the FD interpretation requires assumptions only for the specific difference length $k$, whereas the TWFE interpretation demands these assumptions hold simultaneously across all possible difference lengths. This asymmetry raises a natural question: why aggregate across all available FD estimators to form a TWFE estimator, when each individual FD estimator may provide valid causal identification under weaker conditions?
In traditional panel data econometrics, the answer centers on efficiency gains. Within the textbook fixed effects model, TWFE and all $k$-period FD estimators identify the same population parameter, and the TWFE estimator achieves greater efficiency under additional distributional assumptions about the error term. Indeed, adding a constant effects assumption to the assumptions in Section 3.1 yields this equivalence:
Assumption CE.
(Constant Effects) The per-unit treatment effect defined in Theorem 4 satisfies $\tau_{it}=\tau$ for all $(i,t)$.
Corollary 1.
If Assumptions PO, CT, V, LH, and CE hold, then $\beta_{FE}^{D}=\tau=\beta_{\text{FD},k}^{D}$ for all $k=1,\ldots,T-1$.
However, this efficiency justification becomes irrelevant when the estimators identify different parameters. The possibility of heterogeneous treatment effects immediately undermines this foundation, rendering efficiency comparisons meaningless even when all other identifying assumptions hold.
Moreover, potential bias from violated common trends assumptions at longer horizons can be too substantial to justify the gains from reduced standard errors. Long differences may be less reliable for causal identification if unobserved heterogeneity in trends grows over time. In addition, as noted in Section 3.1, the common trends assumption essentially rules out feedback from past outcome changes to future treatment changes within its time frame. Over longer horizons, this assumption becomes harder to justify, as the likelihood of such feedback increases.
One might alternatively justify TWFE aggregation by appealing to interest in long-run effects, since TWFE regressions exploit long-run differences as well as short-run differences. Identifying effects over $k$-period horizons inevitably requires assuming common trends over at least $k$ periods, making the assumption a necessary compromise on this ground. However, even $k$-period FD coefficients have little connection to $k$-period treatment effects when dynamic effects are present, as formalized in Appendix B.1 under an assumption of homogeneous treatment effects. The TWFE coefficient would therefore represent a difficult-to-interpret weighted average of these already complex effect measures, even under the strong homogeneous effect assumption.
These considerations suggest that TWFE maintains clear advantages only under three joint conditions: (1) common trends assumptions hold for all difference lengths $k$, (2) treatment effects are static rather than dynamic, and (3) treatment effects are linear and homogeneous across units and time periods. When any of these conditions fails, the case for TWFE becomes weaker. Researchers may then focus on FD regressions with specific $k$, which inherit some limitations of TWFE but allow more targeted identifying assumptions. Alternatively, they may employ heterogeneity-robust estimators tailored to their specific setting (e.g., de Chaisemartin and D’Haultfœuille 2020, 2024; Callaway and Sant’Anna 2021; Wooldridge 2021; de Chaisemartin et al. 2022; Borusyak et al. 2024; Callaway et al. 2024), though such methods may be more challenging to apply in broader settings with nonbinary, nonstaggered treatments and time-varying covariates, and some of them still require common trends assumptions across the entire panel duration.
The relationship between TWFE and FD estimators provides valuable diagnostic information about assumption violations. If TWFE and $k$-period FD estimates differ substantially, this signals violation of at least one of the three justifying conditions. Two of these—violations of common trends assumptions for specific difference lengths and the presence of dynamic effects—can be assessed through common trends diagnostics developed in Section 4.2. For the third condition concerning treatment effect homogeneity, the weight diagnostics established in the recent literature (e.g., de Chaisemartin and D’Haultfœuille, 2020) are useful for evaluating how severely violations affect the TWFE coefficient, though with important caveats highlighted in Appendix B.2 that arise from assuming common trends at only one treatment level.
4.2 Diagnosing Violations of the Common Trends Assumption
The common trends assumption cannot be directly tested, since potential outcome changes $\Delta_{k}Y_{it}(d^{0})$ are unobserved. In staggered adoption settings, pre-trend tests provide suggestive evidence about the plausibility of common trends. By contrast, under general treatment paths there is no obvious way to scrutinize this assumption. Yet the plausibility of this assumption can be assessed through testable implications involving the relationship among nonconcurrent treatment and outcome changes. This section develops diagnostic procedures that exploit the temporal structure of panel data to flag violations of common trends across different time horizons.
Sufficient Conditions for Common Trends
The key insight underlying these diagnostics is that the common trends assumption becomes implausible when treatment changes are systematically related to nonconcurrent outcome changes. Suppose that for given $k\in\{2,\ldots,T-1\}$, potential outcome changes in the first $\ell$ periods are correlated with treatment changes in the subsequent $k$-$\ell$ periods, or treatment changes in the first $\ell$ periods are correlated with potential outcome changes in the subsequent $k$-$\ell$ periods. This would suggest intertemporal outcome-to-treatment feedback or unobserved factors that drive both treatment and outcome dynamics over the horizon of $k$ periods, indicating a potential violation of the common trends assumption.
More formally, for each $\ell\in\{1,\ldots,k-1\}$, the following two conditions jointly comprise a sufficient condition for Assumption CT–k:
(Condition I)
$E\left[\Delta_{\ell}Y_{it}(d^{0})|\Delta_{\ell}D_{it},\Delta_{k-\ell}D_{i,t+\ell},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{\ell}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]$ for $t\in\{1,T-k\}$.
(Condition II)
$E\left[\Delta_{k-\ell}Y_{i,t+\ell}(d^{0})|\Delta_{\ell}D_{it},\Delta_{k-\ell}D_{i,t+\ell},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k-\ell}Y_{i,t+\ell}(d^{0})|\Delta_{k}W_{it},C_{i}\right]$ for $t\in\{1,\ldots,T-k\}$.
Condition I requires that potential outcome changes over periods $t$ to $t+\ell$ be mean independent of future treatment changes over periods $t+\ell$ to $t+k$ as well as concurrent treatment changes, conditional on covariates. Condition II requires that potential outcome changes over periods $t+\ell$ to $t+k$ be mean independent of past treatment changes over periods $t$ to $t+\ell$ as well as concurrent treatment changes, conditional on covariates.
Testable Implications and Diagnostic Regressions
To overcome the challenge that potential outcome changes are unobserved, I consider the following decomposition of observed outcome changes, which follows from Assumption PO and the definition of per-unit treatment effects $\tau_{it}$.
$\Delta_{k}Y_{it}=\Delta_{k}Y_{it}(d^{0})+\left(\frac{\tau_{i,t+k}+\tau_{it}}{2}\right)\Delta_{k}D_{it}+\Delta_{k}\tau_{it}\left(\frac{D_{i,t+k}+D_{it}}{2}-d^{0}\right).$
While heterogeneity of per-unit treatment effects $\tau_{it}$ complicates the analysis in general, combining Assumption CE and Condition I yields:
$E\left[\Delta_{\ell}Y_{it}|\Delta_{\ell}D_{it},\Delta_{k-\ell}D_{i,t+\ell},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{\ell}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]+\tau\Delta_{\ell}D_{it}.$
Similarly, under Assumption CE and Condition II:
$E\left[\Delta_{k-\ell}Y_{i,t+\ell}|\Delta_{\ell}D_{it},\Delta_{k-\ell}D_{i,t+\ell},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k-\ell}Y_{i,t+\ell}(d^{0})|\Delta_{k}W_{it},C_{i}\right]+\tau\Delta_{k-\ell}D_{i,t+\ell}.$
These expressions provide the foundation for empirical diagnostics: under the joint hypothesis of Conditions I and II (sufficient for common trends) and constant effects, regressions of outcome changes on concurrent treatment changes and lead/lag treatment changes controlling for covariates ($\Delta_{k}W_{it},C_{i}$) should yield zero coefficients on the lead/lag terms.
The diagnostic procedures can be implemented through the following regression specifications:
$\Delta_{\ell}Y_{it}$ $=$ $\beta_{0}+\beta_{1}\Delta_{k-\ell}D_{i,t+\ell}+\beta_{2}\Delta_{\ell}D_{it}+C_{i}^{\prime}\gamma_{t}+\Delta_{k}W_{it}^{\prime}\delta+\epsilon_{it},$ (14)
$\Delta_{k-\ell}Y_{i,t+\ell}$ $=$ $\beta_{0}+\beta_{1}\Delta_{\ell}D_{it}+\beta_{2}\Delta_{k-\ell}D_{i,t+\ell}+C_{i}^{\prime}\gamma_{t}+\Delta_{k}W_{it}^{\prime}\delta+\epsilon_{it}.$ (15)
Under the joint hypothesis of Conditions I and II (sufficient for Assumption CT–k) and constant effects (Assumption CE), we should observe $\beta_{1}=0$ for each regression. These diagnostics can be implemented for increasing values of $k$, starting with $k=2$. Since Conditions I and II also demand exogeneity of concurrent treatment changes, violation of common trends for some $k^{*}$ implies violation of Conditions I and II for larger $k>k^{*}$, which in turn signals (though does not imply) violation of Assumption CT–k.
These diagnostics resemble distributed-lag regressions but differ in a key respect: these diagnostics exploit treatment changes over nonoverlapping time periods, whereas distributed-lag regressions, being a form of multivariate TWFE regression, pool all possible FDs and therefore incorporate numerous overlapping treatment changes, obscuring the connection to specific common trends assumptions.
Limitations and Interpretation
These diagnostic procedures should not be interpreted as direct tests of the common trends assumption. The link between diagnostic results and the common trends assumption involves two layers of inference, each with important limitations.
First, Conditions I and II are sufficient but not necessary for the common trends assumption (Assumption CT–k). Their violation therefore does not in itself imply that the assumption fails.
Second, $\beta_{1}=0$ in the lead and lag diagnostics is neither sufficient nor necessary for Conditions I and II to hold. It is not sufficient because the diagnostics are silent about the exogeneity of concurrent treatment changes, which is a key part of these conditions. It is not necessary because the diagnostics require the constant effects assumption (Assumption CE). When treatment effects $\tau_{it}$ vary and depend on treatment paths, the coefficients $\beta_{1}$ can be contaminated by this heterogeneity rather than purely reflecting violations of the sufficient conditions.
Despite these limitations, the diagnostics can still be informative for assessing the plausibility of common trends at different horizons, especially since intertemporal outcome–treatment correlation captures economically meaningful threats to identification, such as anticipation, policy responses to past outcomes, or persistent unobserved shocks. They also have the ancillary benefit of detecting dynamic treatment effects, since $\beta_{1}\neq 0$ can arise when treatment effects depend on treatment history (Appendix B.1). Accordingly, the diagnostics help evaluate whether TWFE estimates are likely to be interpretable as causal effects, and whether estimators that rely more heavily on short-run (or long-run) variation may be preferable.
5 Empirical Illustration
This section illustrates insights from the econometric results by studying the impact of minimum wages on employment using TWFE regressions. The TWFE regression analysis based on state-level panel data in the US has attracted considerable attention for decades. The econometric framework of this paper, applicable to a broad class of multiperiod panel settings, is particularly well-suited for this empirical context. The treatment variable, the state minimum wage, is continuous and changes many times in a given state. The existing insights from the TWFE regression literature, typically addressing binary treatments and staggered adoption designs, do not directly apply to this setting.
To present the estimates, I use a state–year panel of employment outcomes and minimum wage laws in 50 US states and the District of Columbia from 1979 to 2019. As a dependent variable in a TWFE regression, I use the log employment rate of teens (ages 16–19) from the Current Population Survey (CPS). Another dependent variable is the net job creation rate from the Business Dynamics Statistics (BDS), using four major sectors with high shares of workers without college education. [^10] This sample choice follows Manning (2021), who uses CPS data from 1979 to 2019, aggregated into quarterly state-level observations. However, my analysis is based on annual frequency to incorporate the BDS data. The four sectors from the BDS data that I focus on are: Construction, Manufacturing, Retail Trade, and Accommodation and Food Services. Using the data from all sectors does not qualitatively change the estimates presented below, but the estimated effects become smaller in magnitude. These employment measures, as stock and flow variables, capture different aspects of labor market adjustment. The focus on age groups or sectors more likely affected by the minimum wage is practical, since the majority of workers in the US are not directly affected by the minimum wage. The minimum wage data are from Vaghul and Zipperer (2019), and I use the log of the annual average of the state minimum wage as the treatment variable in each regression. I use census region indicators as covariates.
I begin by illustrating the weighted-average characterization suggested in Section 2.2. Figure 1 presents the FD coefficients $\widehat{\beta}_{\text{FD},k}$ for $k=1,\ldots,40$ as dots accompanied by standard error bars, the TWFE coefficient $\widehat{\beta}_{\text{FE}}$ as a horizontal dashed line, and the weights given by Theorem 3 as bars. Since all covariates are time-invariant, the adjustment term in Theorem 3 can be ignored. While the inclusion of time-varying covariates would result in a deviation from the exact weighted-average interpretation, Appendix E.1 finds that the deviation is minimal in this setting even after including time-varying covariates such as log per-capita income and log teen population.
Figure 1: The TWFE and FD Coefficients on Log Minimum Wage
(a) Log Employment Rate of Teens
(b) Net Job Creation Rate (%)
  • Notes: The FD coefficients are illustrated with standard error bars. Covariates are census region indicators. The standard errors are robust to to heteroskedasticity and correlation across observations on the same state.
In Panel (a) of Figure 1, the TWFE coefficient is $\widehat{\beta}_{\text{FE}}=-0.266$, whereas the FD coefficient is $\widehat{\beta}_{\text{FD},1}=-0.001$ with a one-year gap and becomes larger in magnitude as the gap increases. While the difference among “short” FD, “long” FD, and TWFE estimates itself has long been recognized in the minimum wage effects literature (e.g., Neumark and Wascher, 1992), an explicit numerical relationship among these estimates has not been previously established. The theoretical results in Section 2 demonstrate that a TWFE regression aggregates these heterogeneous FD coefficients into a single coefficient. The weighted average of all 40 FD coefficients is indeed 0.266-0.266. Panel (b) of Figure 1 presents analogous plots for a regression of the net job creation rate. Even though the TWFE coefficient is $\widehat{\beta}_{\text{FE}}=-2.65$, one-year FD coefficient $\widehat{\beta}_{\text{FD},1}=-5.74$ is larger in magnitude and some long-run FD coefficients have different signs. The weighted average of the FD coefficients is again confirmed to be identical to the TWFE coefficient.
The weights depicted in Figure 1 are identical across the two panels. This consistency arises because the weights depend only on the explanatory variables, not on the dependent variables. As observed in (10), these weights are associated with the unexplained variation of treatment changes. Consequently, the weights exhibit a distinct pattern: they are small for both the short- and long-run FD coefficients, albeit for different reasons. In the case of the short-run FD coefficients, the weights are small because minimum wage adjustments tend to be incremental over short periods. In contrast, for the long-run FD coefficients, the weights are small due to the limited number of observations of long-run changes.
The substantial differences among FD coefficients documented in Figure 1 raise fundamental questions about causal interpretation. Traditional panel data econometrics justifies TWFE’s aggregation of these estimates on efficiency grounds, but this efficiency justification is valid only when all FD estimators identify the same population parameter. The patterns observed here suggest that different FD estimators likely identify distinct parameters rather than noisy estimates of a common effect. Indeed, the hypothesis that these coefficients are all identical can be rejected even at a significance level of 0.01%. More concerning, some of these estimates may not identify causal parameters at all if the common trends assumption is violated. With a 41-year panel spanning 1979–2019, it seems unlikely that this assumption holds uniformly across all k=1,,40k=1,...,40. The procedures developed in Section 4.2 provide a systematic way to evaluate these concerns across different time horizons.
Table 1 presents regressions based on equations (14) and (15). These diagnostics test whether past and future treatment changes have zero coefficients when regressing outcome changes on them, controlling for concurrent treatment changes and covariates. For practical presentation, I focus on even values of $k\leq 10$ with lead/lag length $\ell=k/2$.
Table 1: Diagnosing Common Trends Violation for Minimum Wage Changes
$k=2$ $k=4$ $k=6$ $k=8$ $k=10$
Lead Lag Lead Lag Lead Lag Lead Lag Lead Lag
Panel A. Log Teen Employment Rate
Future change 0.101 0.081 –0.013 –0.094 –0.092
(0.061) (0.056) (0.056) (0.061) (0.050)
Past change –0.128 0.017 0.024 –0.021 –0.121
(0.070) (0.040) (0.076) (0.085) (0.116)
Concurrent –0.034 0.027 0.025 –0.008 0.032 0.036 –0.002 0.052 0.009 0.013
change (0.061) (0.058) (0.067) (0.066) (0.092) (0.067) (0.107) (0.062) (0.118) (0.054)
P-value 0.019 0.288 0.936 0.048 0.002
Panel B. Net Job Creation Rate (%)
Future change 1.88 4.24 2.99 2.79 4.18
(1.23) (1.53) (1.16) (1.28) (1.48)
Past change 0.19 –4.24 –3.58 –1.45 –1.32
(2.60) (1.55) (1.53) (2.12) (2.27)
Concurrent –6.17 –5.98 –4.48 –5.10 –5.42 –5.55 –4.87 –5.00 –4.13 –5.16
change (2.35) (2.72) (1.78) (1.64) (1.87) (1.46) (1.88) (1.22) (2.16) (1.31)
P–value 0.280 0.004 0.017 0.037 0.005
  • Note: Each column reports a regression of outcome changes on future or past minimum wage changes using equation (14) or (15) with $\ell=k/2$, controlling for concurrent minimum wage changes and census region indicators. Standard errors are in parentheses and robust to heteroskedasticity and correlation across observations on the same state. P–values test the joint null that future and past change coefficients equal zero.
For teen employment, the hypothesis that both past and future treatment change coefficients equal zero is rejected at the 5% significance level for $k=$ 2, 8, and 10. For $k=2$, the future change coefficient (0.101) and the past treatment change coefficient (–0.128) exceed the concurrent change coefficient in magnitude, suggesting short-term feedback mechanisms or dynamic treatment effects. For $k=$ 8 and 10, both past and future minimum wage changes have negative coefficients, which raises the possibility of unobserved long-term factors that simultaneously drive minimum wage policies and teen employment trends over extended periods.
The job creation rate estimates also exhibit patterns that are hard to reconcile with the common trends assumption. The hypothesis of zero coefficients is rejected at the 5% level for $k=$ 4, 6, 8, and 10. Most notably, the future treatment change coefficients are consistently positive and significant across these horizons, indicating that increases in net job creation systematically precede minimum wage increases. This pattern may reflect intertemporal outcome-to-treatment feedback, where policymakers respond to improving labor market conditions by raising minimum wages.
Overall, the diagnostics indicate that the common trends assumption is most credible only at short horizons. This suggests that estimators relying exclusively on short-run changes may be more appropriate than the TWFE estimator for identifying causal effects in this setting. Yet, because employment is a stock variable that may adjust only gradually, an exclusive focus on short-run variation may never recover the economically relevant long-run effect, as emphasized by Baker et al. (1999), Sorkin (2015), and Meer and West (2016), among many others. This tension underscores the difficulty of designing empirical strategies that are both credible and substantively informative, and highlights the value of diagnostics in making such trade-offs explicit.
6 Conclusion
Building on a long-overlooked numerical equivalence between TWFE and pooled FD regressions, this paper provides insights into how TWFE regressions translate data into the coefficient of interest. A key observation is that TWFE regressions reflect associations among changes in the variables of interest, pooling all possible short- and long-run comparisons. As a natural consequence of this property, causal interpretation of the TWFE coefficient requires the common trends assumption to hold across all time horizons. To address this challenge, I develop diagnostic procedures that exploit testable implications to help researchers assess whether the common trends assumptions are plausible across different time horizons.
This paper focuses on diagnosing problems of TWFE regressions rather than proposing solutions. While recent literature has developed various alternative estimators in specific contexts such as binary treatments or staggered adoption designs, extending robust causal inference methods to broader settings with general treatment paths and time-varying covariates remains an important challenge for future research.
Appendix A Proofs
As defined in the main paper, $\Delta_{k}a_{t}\equiv a_{t+k}-a_{t}$ denotes a $k$-period change, and $\overline{a}\equiv\frac{1}{T}\sum_{t=1}^{T}a_{t}$ represents the average of any time series $\{a_{t}\}_{t=1}^{T}$. Some proofs use the following identify, which follows from basic algebra and is a well-known property of U-statistics regarding the sample covariance. For any two real sequences $\{a_{t}\}_{t=1}^{T}$ and $\{b_{t}\}_{t=1}^{T}$,
$\frac{1}{T}\sum_{t=1}^{T}\left(a_{t}-\overline{a}\right)\left(b_{t}-\overline{b}\right)=\frac{1}{T^{2}}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\Delta_{k}a_{t}\Delta_{k}b_{t}.$ (A.1)
The following algebraic identify will also be used by some proofs. Let $A(s,t)$ be any real-valued function that satisfies A(s,t)=A(t,s)A(s,t)=-A(t,s) and $A(t,t)=0$ and let $\{b_{t}\}_{t=1}^{T}$ be any real sequence. Then,
$\sum_{s>t}\left(b_{s}-b_{t}\right)A(s,t)=\sum_{t>s}b_{t}A(t,s)+\sum_{s>t}b_{t}A(t,s)=\sum_{t=1}^{T}\sum_{s=1}^{T}b_{t}A(t,s).$ (A.2)
Proof of Theorem 1
Defining $e_{it}\equiv Y_{it}-X_{it}^{\prime}\beta-\alpha_{i}-C_{i}^{\prime}\gamma_{t}-\mu_{t}$ and using (A.1), the TWFE objective (4) can be rewritten as
$\sum_{i=1}^{N}\sum_{t=1}^{T}e_{it}^{2}=\frac{1}{T}\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left(\Delta_{k}e_{it}\right)^{2}+T\sum_{i=1}^{N}\overline{e}_{i}^{2}.$ (A.3)
The first term is identical to (5). For any $\left(\beta,\{\gamma_{t},\mu_{t}\}_{t=1}^{T}\right)$, letting $\alpha_{i}=\overline{Y}_{i}-\overline{X}_{i}^{\prime}\beta-C_{i}^{\prime}\overline{\gamma}-\overline{\mu}$ results in $\overline{e}_{i}=0$ for each $i=1,\ldots,N$. Therefore, $\left(\beta,\{\gamma_{t},\mu_{t}\}_{t=1}^{T}\right)$ minimizing (5) constitutes the solution to (4).
Proof of Theorem 2
The FWL theorem implies that each $\widehat{\beta}_{\text{FD},k}$ is given by
$\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right)\widehat{\beta}_{\text{FD},k}$ $=$ $\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{Y}_{it}\right),$ (A.4)
where $\widehat{\beta}_{\text{FD},k}$ is uniquely defined if and only if $\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}$ is positive definite.
Solving the least-squares problem (4), $\widehat{\beta}_{\text{FE}}$ is given by
$\sum_{i=1}^{N}\sum_{t=1}^{T}\left(\widetilde{X}_{it}-\frac{1}{T}\sum_{s=1}^{T}\widetilde{X}_{is}\right)\left(\widetilde{X}_{it}-\frac{1}{T}\sum_{s=1}^{T}\widetilde{X}_{is}\right)^{\prime}\widehat{\beta}_{\text{FE}}\\ =\sum_{i=1}^{N}\sum_{t=1}^{T}\left(\widetilde{X}_{it}-\frac{1}{T}\sum_{s=1}^{T}\widetilde{X}_{is}\right)\left(\widetilde{Y}_{it}-\frac{1}{T}\sum_{s=1}^{T}\widetilde{Y}_{is}\right).$ (A.5)
Applying (A.1) to each element of matrices in (A.5),
$\widehat{\beta}_{\text{FE}}=\left(\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right)^{-1}\left(\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{Y}_{it}\right),$ (A.6)
provided that $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}$ is positive definite.
Note that (A.6) is a solution not only to (4) but also to
$\underset{\beta,\{\gamma_{k,t}^{*},\mu_{k,t}^{*}\}_{k,t}}{\min}\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{k,t}^{*}-\mu_{k,t}^{*}\right)^{2},$ (A.7)
given the linearity of the residualizing transformation (7). Imposing the separability restrictions, $\gamma_{k,t}^{*}=\gamma_{t+k}-\gamma_{t}$ and $\mu_{k,t}^{*}=\mu_{t+k}-\mu_{t}$, on this regression results in (5). However, these separability restrictions do not bind, since the two least-squares problems provide the same solution $\widehat{\beta}_{\text{FE}}$.
Combining (A.4) and (A.6), the TWFE estimator can be expressed as a matrix-weighted average $\widehat{\beta}_{\text{FE}}=\sum_{k=1}^{T-1}\widehat{\varOmega}_{k}\widehat{\beta}_{\text{FD},k}$ with
$\widehat{\varOmega}_{k}=\left(\sum_{\ell=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-\ell}\Delta_{\ell}\widetilde{X}_{it}\Delta_{\ell}\widetilde{X}_{it}^{\prime}\right)^{-1}\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{X}_{it}\Delta_{k}\widetilde{X}_{it}^{\prime}\right),$
where the weight matrices $\widehat{\varOmega}_{k}$ sum to the identity matrix, and are all positive definite if and only if every $\widehat{\beta}_{\text{FD},k}$ is uniquely defined.
Proof of Theorem 3
A basic regression algebra yields
$\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)$
$=$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}Y_{it}\left\{\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FD},k}^{W}\right)+\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)\right\}$
$=$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left\{\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FD},k}^{W}\right)+\Delta_{k}Y_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)\right\}$
$=$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left\{\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)+\left(\Delta_{k}Y_{it}-\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\right)\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)\right\}$
$=$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left\{\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)+\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)^{\prime}\left(\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right)\widehat{\beta}_{\text{FD},k}^{W}\right\}$
$=$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\left\{\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FE}}^{W}\right)+\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)^{\prime}\left(\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right)\left(\widehat{\beta}_{\text{FD},k}^{W}-\widehat{\beta}_{\text{FE}}^{W}\right)\right\},$
where the second equality follows from
$\widehat{\beta}_{\text{FD},k}^{D}=\frac{\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FD},k}^{W}\right)}{\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\widehat{\delta}_{\text{FD},k}^{W}\right)},$
the fourth equality uses
$\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right)\widehat{\beta}_{\text{FD},k}^{W}=\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{W}_{it}\left(\Delta_{k}Y_{it}-\widehat{\beta}_{\text{FD},k}^{D}\Delta_{k}D_{it}\right),$
and the last equality uses
$\sum_{k=1}^{T-1}\left(\widehat{\delta}_{\text{FD},k}^{W}-\widehat{\delta}_{\text{FE}}^{W}\right)^{\prime}\left(\sum_{i=1}^{N}\sum_{t=1}^{T-k}\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right)=0,$
which follows from Theorem 2. The other equalities merely involve addition and subtraction.
Proof of Lemma 1
Assumption LH implies
$E\left[\Delta_{k}\widetilde{D}_{it}|\Delta_{k}W_{it},C_{i}\right]=\Delta_{k}\widetilde{W}_{it}^{\prime}\delta^{W}.$
Then, the FD coefficients for $k=1,\ldots,T-1$ are given by
$\delta_{\text{FD},k}^{W}$ $=\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{D}_{it}\right]\right)$
$=\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}E\left[\Delta_{k}\widetilde{D}_{it}|\Delta_{k}W_{it},C_{i}\right]\right]\right)$
$=\delta^{W}.$
In addition, the TWFE coefficient is given by
$\delta_{\text{FE}}^{W}$ $=\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{D}_{it}\right]\right)$
$=\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}E\left[\Delta_{k}\widetilde{D}_{it}|\Delta_{k}W_{it},C_{i}\right]\right]\right)$
$=\delta^{W}.$
Proof of Theorem 4
Using Assumption L, the denominator in (11) is given by
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]\right)^{2}+\left\{\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\delta_{k,t}^{W}-\delta_{\text{FE}}^{W}\right)\right\}^{2}\right],$
which is strictly positive under Assumption V.
Using identity (A.2), the denominator in (11) can be alternatively expressed as
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{s>t}E\left[\left((D_{is}-d^{0})-(D_{it}-d^{0})\right)\left(\widetilde{D}_{is}-\widetilde{D}_{it}-\left(\widetilde{W}_{is}-\widetilde{W}_{it}\right)^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{t=1}^{T}E\left[(D_{it}-d^{0})\sum_{s=1}^{T}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{\text{FE}}^{W}\right)\right].$ (A.8)
Now, using the definition of $\tau_{it}$, decompose observed outcome as $Y_{it}=\tau_{it}(D_{it}-d^{0})+Y_{it}(d^{0})$ and plug it into the numerator in (11). The treatment effect component yields
$\sum_{s>t}E\left[\left(\tau_{is}(D_{is}-d^{0})-\tau_{it}(D_{it}-d^{0})\right)\left(\widetilde{D}_{is}-\widetilde{D}_{it}-\left(\widetilde{W}_{is}-\widetilde{W}_{it}\right)^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{t=1}^{T}E\left[\tau_{it}(D_{it}-d^{0})\sum_{s=1}^{T}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{\text{FE}}^{W}\right)\right].$ (A.9)
In addition, the baseline outcome component yields
$E\left[\Delta_{k}Y_{it}(d^{0})\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]$ $=E\left[E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]$
$=E\left[E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]$
$=E\left[\Delta_{k}Y_{it}(d^{0})\left(E\left[\Delta_{k}\widetilde{D}_{it}|\Delta_{k}W_{it},C_{i}\right]-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]$
$=E\left[\Delta_{k}Y_{it}(d^{0})\Delta_{k}\widetilde{W}_{it}^{\prime}\left(\delta_{k,t}^{W}-\delta_{\text{FE}}^{W}\right)\right].$ (A.10)
The first and third equalities use the law of iterated expectations. The second equality uses Assumption CT, and the fourth equality uses Assumption L.
Combining (A.8), (A.9) and (A.10) yields $\beta_{\text{FE}}^{D}=\sum_{t=1}^{T}E\left[\omega_{it}\tau_{it}\right]+B$, where the weights $\omega_{it}$ satisfy $\sum_{t=1}^{T}E[\omega_{it}]=1$ and
$\omega_{it}$ $\propto$ $(D_{it}-d^{0})\sum_{s=1}^{T}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{\text{FE}}^{W}\right)$
$\propto$ $(D_{it}-d^{0})\left(\widetilde{D}_{it}-\overline{\widetilde{D}}_{i}-\left(\widetilde{W}_{it}-\overline{\widetilde{W}}_{i}\right)^{\prime}\delta_{\text{FE}}^{W}\right).$
Proof of Theorem 5
The population FD coefficient on $\Delta_{k}D_{it}$ is given by
$\beta_{\text{FD},k}^{D}=\frac{\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]}{\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]},$ (A.11)
where
$\delta_{\text{FD},k}^{W}=\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{D}_{it}\right]\right)$
represents the population version of $\widehat{\delta}_{\text{FD},k}^{W}$.
Using law of iterated expectations and Assumption LH–k,
$\delta_{\text{FD},k}^{W}$ $=\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}\Delta_{k}\widetilde{W}_{it}^{\prime}\right]\right)^{-1}\left(\sum_{t=1}^{T-k}E\left[\Delta_{k}\widetilde{W}_{it}E\left[\Delta_{k}\widetilde{D}_{it}|\Delta_{k}\widetilde{W}_{it}\right]\right]\right)$
$=\delta_{k}^{W},$
and
$\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}=\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right].$ (A.12)
Then, the denominator in (A.11) is given by
$\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]=\sum_{t=1}^{T-k}E\left[\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]\right)^{2}\right],$
which is strictly positive under Assumption V.
Using identity (A.2), the denominator in (A.11) can be alternatively expressed as
$\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{s>t}\mathbbm{1}_{|t-s|=k}E\left[\left((D_{is}-d^{0})-(D_{it}-d^{0})\right)\left(\widetilde{D}_{is}-\widetilde{D}_{it}-\left(\widetilde{W}_{is}-\widetilde{W}_{it}\right)^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{t=1}^{T}E\left[(D_{it}-d^{0})\sum_{s=1}^{T}\mathbbm{1}_{|t-s|=k}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{k}^{W}\right)\right].$ (A.13)
Using identity (A.2),
$\sum_{s>t}\mathbbm{1}_{|t-s|=k}E\left[\left(\tau_{is}(D_{is}-d^{0})-\tau_{it}(D_{it}-d^{0})\right)\left(\widetilde{D}_{is}-\widetilde{D}_{it}-\left(\widetilde{W}_{is}-\widetilde{W}_{it}\right)^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{t=1}^{T}E\left[\tau_{it}(D_{it}-d^{0})\sum_{s=1}^{T}\mathbbm{1}_{|t-s|=k}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{k}^{W}\right)\right].$ (A.14)
In addition,
$E\left[\Delta_{k}Y_{it}(d^{0})\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]$ $=E\left[E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]$
$=E\left[E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD,k}}^{W}\right)\right]$
$=E\left[E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i}\right]\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]\right)\right]$
$=0$ (A.15)
The first and fourth equalities use the law of iterated expectations. The second equality uses Assumption CT, and the third equality uses (A.12).
Plugging $Y_{it}=\tau_{it}(D_{it}-d^{0})+Y_{it}(d^{0})$ into the numerator in (A.11) and combining (A.14) and (A.15) yields
$\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{t=1}^{T}E\left[\tau_{it}(D_{it}-d^{0})\sum_{s=1}^{T}\mathbbm{1}_{|t-s|=k}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{k}^{W}\right)\right].$ (A.16)
Combining (A.13) and (A.16), the weights $\omega_{it}^{(k)}$ satisfy $\sum_{t=1}^{T}E[\omega_{it}^{(k)}]=1$ and
$\omega_{it}^{(k)}$ $\propto$ $(D_{it}-d^{0})\sum_{s=1}^{T}\mathbbm{1}_{|t-s|=k}\left(\widetilde{D}_{it}-\widetilde{D}_{is}-\left(\widetilde{W}_{it}-\widetilde{W}_{is}\right)^{\prime}\delta_{k}^{W}\right).$
Appendix B Key Extensions
This section extends Section 3 in two key directions. Proofs of additional theoretical results are provided at the end of this section.
B.1 Dynamic Effects
The framework in Section 3 restricts potential outcomes to depend only on current treatment status, ruling out dynamic effects where past treatments influence current outcomes. This section relaxes this restriction to allow for dynamic treatment effects, though under a constant effects assumption to maintain tractability. While this assumption is restrictive, the analysis serves to illustrate that even under this strong assumption, $k$-period FD coefficients lack clear interpretation as $k$-period treatment effects, and the TWFE coefficient is even more difficult to interpret.
Assumption PO–D.
(Potential Outcome, Dynamic) For each $t=1,\ldots,T$, $\{Y_{it}(d_{t},\ldots,d_{t-L}):(d_{t}\ldots,d_{t-L})\in(\underline{d},\overline{d})^{L+1}\}$ is a stochastic process that defines a potential outcome associated with each possible treatment level $(d_{t}\ldots,d_{t-L})\in(\underline{d},\overline{d})^{L+1}$, where $\infty\leq\underline{d}<\overline{d}\leq\infty$. The observed outcome is given by $Y_{it}=Y_{it}(D_{it},\ldots,D_{i,t-L})$.
Assumption CT–D.
(Conditional Common Trends, Dynamic) There exists $d^{0}\in(\underline{d},\overline{d})$ such that
$E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})|\Delta_{k}W_{it},C_{i}\right]$
for any $(k,t)$ with $1\leq t<t+k\leq T$.
Assumption CE–D.
(Constant Effects, Dynamic) The potential outcome function satisfies $Y_{it}(d_{t}^{\prime},\ldots,d_{t-L}^{\prime})-Y_{it}(d_{t},\ldots,d_{t-L})=\sum_{\ell=0}^{L}\gamma_{\ell}(d_{t-\ell}^{\prime}-d_{t-\ell})$ for any $(d_{t}^{\prime}\ldots,d_{t-L}^{\prime}),(d_{t}\ldots,d_{t-L})\in(\underline{d},\overline{d})^{L+1}$.
Assumption PO–D extends the potential outcome framework to allow outcomes to depend on treatment history up to $L$ periods in the past, including the treatment status before the beginning of the panel. This encompasses a broad class of dynamic effects, from immediate impacts to persistent effects that decay or accumulate over time. Assumption CT–D maintains the common trends requirement but now applies to potential outcomes under a constant treatment path at the baseline level $d^{0}$. Assumption CE–D requires that the impact of treatment $\ell$ periods ago be constant across units, time periods, and treatment histories.
Under these assumptions, the parameters $\{\gamma_{\ell}\}_{\ell=0}^{L}$ capture the dynamic structure of treatment effects. The contemporary effect is $\gamma_{0}$, while $\gamma_{\ell}$ for $\ell>0$ represents the additional impact on current outcomes from treatment received $\ell$ periods ago, holding all other treatment history constant. Partial sums of $\{\gamma_{\ell}\}_{\ell=0}^{L}$ represent cumulative treatment effects at different horizons. For example, if the treatment increases by one unit once and for all, its impact on the outcome $K$ periods later is $\sum_{\ell=0}^{\min\{K,L\}}\gamma_{\ell}$.
Theorem B1.
Under Assumptions PO–D, CT–D, V, and LH,
$\beta_{\text{FD},k}^{D}=\gamma_{0}+\sum_{\ell=1}^{L}\frac{\sum_{t=1}^{T-k}E\left[Cov\left(\Delta_{k}D_{it},\Delta_{k}D_{i,t-\ell}|\Delta_{k}W_{it},C_{i}\right)\right]}{\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right]}\gamma_{\ell},$
$\beta_{\text{FE}}^{D}=\gamma_{0}+\sum_{\ell=1}^{L}\frac{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Cov\left(\Delta_{k}D_{it},\Delta_{k}D_{i,t-\ell}|\Delta_{k}W_{it},C_{i}\right)\right]}{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right]}\gamma_{\ell},$
Theorem B1 characterizes how FD and TWFE coefficients relate to the underlying dynamic effect parameters. While both $\beta_{\text{FD},k}^{D}$ and $\beta_{\text{FE}}^{D}$ fully capture the contemporary effect parameter $\gamma_{0}$, the extent to which each coefficient captures dynamic effects $\gamma_{\ell}$ for $\ell>0$ depends on how much current treatment changes $\Delta_{k}D_{it}$ predict lagged treatment changes $\Delta_{k}D_{i,t-\ell}$.
With $\ell=k$, there is no mechanical correlation between $\Delta_{k}D_{it}$ and $\Delta_{k}D_{i,t-k}$ since these represent treatment changes over non-overlapping time periods. While these treatment changes can still be correlated through underlying treatment dynamics, there is no inherent connection between $k$-period FD coefficients and treatment effects of $k$-period horizon, contrary to what might be intuitively expected. The TWFE coefficient becomes even more difficult to interpret, as it represents a complex weighted average across all difference lengths $k$ and lag lengths $\ell$, with weights determined by the full correlation structure of treatment paths across the entire panel.
The presence of dynamic effects affects the diagnostic procedures developed in Section 4.2. Under Assumptions PO–D and CE–D, observed outcome changes decompose as:
$\Delta_{k}Y_{it}=\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})+\sum_{\ell=0}^{L}\gamma_{\ell}\Delta_{k}D_{i,t-\ell}.$
The decomposition shows that outcome changes $\Delta_{k}Y_{it}$ depend on lagged treatment changes $\Delta_{k}D_{i,t-\ell}$ for $\ell>0$. In lag diagnostics, some of these lagged changes directly appear as regressors. Even in lead diagnostics, future treatment changes may correlate with these lagged changes. Therefore, diagnostic procedures in Section 4.2 flag dynamic effects in addition to violations of common trends assumptions.
B.2 No Baseline Treatment Level
Even though a natural baseline treatment level such as “no treatment” exists in some applications, in other applications making Assumption CT only for a particular treatment level $d^{0}$ can be too arbitrary. Moreover, even when a natural baseline exists, there is typically no clear theoretical justification for why common trends should hold exclusively at this level while being allowed to fail at all others, absent dynamic effects. This assumption affects weight diagnostics commonly used to assess TWFE sensitivity to treatment effect heterogeneity. The prominence of negative weights in recent literature depends crucially on assuming common trends at only one level; Fabre (2025) shows that in binary treatment settings with static effects, assuming common trends for both untreated and treated potential outcomes allows the TWFE coefficient to be expressed as a convex combination of treatment effects.
The following analysis addresses these issues by requiring common trends to hold at all treatment levels.
Assumption CTA.
(Conditional Common Trends at Any Level of Treatment) For any $d\in(\underline{d},\overline{d})$ and any $(k,t)$ with $1\leq t<t+k\leq T$,
$E\left[\Delta_{k}Y_{it}(d)|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d)|\Delta_{k}W_{it},C_{i}\right].$
Assumption LC.
(Lipschitz Continuity) For each $t=1,\ldots,T$, there exists a random variable $M_{it}$ that satisfies $|Y_{it}(d^{\prime})-Y_{it}(d)|\leq M_{it}|d^{\prime}-d|$ for any $d^{\prime},d\in(\underline{d},\overline{d})$, $E\left[M_{it}^{2}\right]<\infty$, and $E\left[M_{it}^{2}D_{it}^{2}\right]<\infty$.
Assumption CTA extends Assumption CT to every possible treatment level. Even though Assumption CTA is technically stronger than Assumption CT, it can be practically more reasonable, particularly in settings where no natural baseline treatment level exists. For instance, with minimum wage policy, assuming parallel trends only at \$10/hour but not at \$9 or \$11 lacks theoretical justification. In such cases, Assumption CTA avoids the need to defend why parallel trends should hold exclusively at a single, arbitrarily chosen treatment level while being allowed to fail at all others. An alternative baseline-free approach may assume common trends for the status-quo outcome $\Delta_{k}Y_{it}(D_{it})$ conditional on the same initial treatment status $D_{it}$, akin to a strategy employed by de Chaisemartin et al. (2024). This alternative assumption can support causal interpretation for estimators that control for initial treatment status, but it cannot for the standard TWFE estimator, which does not control for initial treatment status.
Assumption LC is a regularity condition which ensures that a potential outcome process $Y_{it}(d)$ is differentiable almost everywhere and is sufficiently smooth. Under this differentiability, the derivative $Y_{it}^{\prime}(d)$ represents the marginal causal effect of treatment at level dd. [^11] This definition does not preclude discrete treatment settings, since potential outcomes $Y_{it}(d)$ defined for discrete $d$ can be extended to the real line without loss of generality. While my exposition uses derivatives and integrals for continuous treatments, the results apply to discrete treatments by replacing them with differences and summations. Given this definition, Assumption CTA also results in conditional common trends for treatment effects $Y_{it}^{\prime}(d)$, implying that the treatment effects are additive in unit-specific and period-specific terms. This contrasts with Assumption CT, which applies only to a baseline treatment level and allows arbitrary treatment effect heterogeneity.
I maintain Assumptions PO, V, and LH, which are also used in Section 3.1. I focus on a setting under Assumption LH for simplicity; the bias term $B$ does not differ from Theorem 4 when Assumption LH is replaced by Assumption L.
It turns out that these assumptions are sufficient for providing a weighted-average interpretation of a TWFE coefficient, but not sufficient for an unambiguous weighted-average interpretation.
Theorem B2.
Suppose Assumptions PO, V, LH, CTA, and LC hold. Then, for any $d^{0}\in(\underline{d},\overline{d})$ and any $g:(\underline{d},\overline{d})\mapsto\mathbb{R}$ with $\int_{\underline{d}}^{\overline{d}}|g(x)|dx<\infty$,
$\beta_{\text{FE}}^{D}=\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(x)\left(\psi_{it}(x;d^{0})+g(x)R_{it}\right)\right]dx,$
where $R_{it}=\left(\widetilde{D}_{it}-\overline{\widetilde{D}}_{i}\right)-\left(\widetilde{W}_{it}-\overline{\widetilde{W}}_{i}\right)^{\prime}\delta_{\text{FE}}^{W}$ is a TWFE residual and the weight function
$\psi_{it}(d;d^{0})\equiv\frac{\left(\mathbbm{1}_{D_{it}\geq d}-\mathbbm{1}_{d^{0}\geq d}\right)R_{it}}{\sum_{s=1}^{T}E\left[D_{is}R_{is}\right]}$
satisfies $\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[\psi_{it}(x;d^{0})\right]dx=1$.
Theorem B2 expresses the population TWFE coefficient $\beta_{\text{FE}}^{D}$ as a weighted average of treatment effects $Y_{it}^{\prime}(d)$, where weights sum to one but may be negative. However, there are infinitely many weight functions that can yield this weighted-average expression. This multiplicity stems from two sources: the choice of baseline treatment $d^{0}$, and more broadly, the possibility to add any function proportional to the TWFE residual $R_{it}$. This issue arises because Assumption CTA implies that treatment effects $Y_{it}^{\prime}(d)$ are additive in unit and period effects. Due to this assumption, $Y_{it}^{\prime}(d)$ is orthogonal to the TWFE residual $R_{it}$, that is, $\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(d)R_{it}\right]=0$ for any $d\in(\underline{d},\overline{d})$. As a result, any function proportional to $R_{it}$, such as $\mathbbm{1}_{d^{0}\geq d}R_{it}$, can be added to or subtracted from the weight function without affecting the weighted average.
This property poses challenges beyond ambiguity in the weighted-average interpretation: it can also exaggerate the sensitivity of the TWFE coefficient to treatment effect heterogeneity. This sensitivity is often assessed through the variance of weights or the proportion of negative weights (e.g., de Chaisemartin and D’Haultfœuille, 2020). While these measures remain valid when common trends hold only for a baseline treatment level, they inflate under Assumption CTA due to extraneous variation unrelated to treatment effects, making the TWFE coefficient appear more sensitive to heterogeneity than it may actually be.
To address this issue, I introduce additional conditions that characterize the exogeneity requirements more explicitly than Assumption CTA. These conditions help eliminate the weight variation orthogonal to $Y_{it}^{\prime}(d)$, yielding a more interpretable weighted-average representation.
Assumption SE.
(Strict Exogeneity) For any $d\in(\underline{d},\overline{d})$ and $(k,t)$ with $1\leq t<t+k\leq T$,
$E\left[\Delta_{k}Y_{it}(d)|D_{i,t+k},D_{it},\overline{D}_{i},W_{i,t+k},W_{it},\overline{W}_{i},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d)|W_{i,t+k},W_{it},\overline{W}_{i},C_{i}\right].$
Assumption S.
(Sufficiency) For any $d\in(\underline{d},\overline{d})$ and $(k,t)$ with $1\leq t<t+k\leq T$,
$E\left[\Delta_{k}Y_{it}(d)|W_{i,t+k},W_{it},\overline{W}_{i},C_{i}\right]=E\left[\Delta_{k}Y_{it}(d)|\Delta_{k}W_{it},C_{i}\right].$
Assumption SE is a nonparametric and weaker version of a strict exogeneity condition that is common in panel data models. Unlike traditional strict exogeneity, which typically requires the entire treatment series $(D_{i1},\ldots,D_{iT})$ to be exogenous—as seen in condition (13) used by de Chaisemartin and D’Haultfœuille (2020)—Assumption SE requires exogeneity of only the pair $(D_{i,t+k},D_{it})$ and the time-series average $\overline{D}_{i}$. Similarly, this assumption conditions only on $\left(W_{i,t+k},W_{it},\overline{W}_{i}\right)$ for time-varying covariates, avoiding reliance on the full covariate history. Nevertheless, the time-series treatment average $\overline{D}_{i}$ depends on past and future treatment statuses beyond the specified time frame $[t,t+k]$. As noted in Section 3.1, Assumption CT (and CTA) collectively impose an exogeneity requirement that rules out the possibility that potential outcome changes $\Delta_{k}Y_{it}(d)$ influence future treatment assignments or correlate with past treatment assignments, mitigating concerns about this constraint. Including $\overline{D}_{i}$ and $\overline{W}_{i}$ in the time-invariant covariates $C_{i}$, though uncommon in practice, would further relax these requirements.
Assumption S ensures that concurrent covariate changes $\Delta_{k}W_{it}$ are sufficient for predicting potential outcome changes $\Delta_{k}Y_{it}(d)$. This assumption is necessary because Assumption SE alone does not imply Assumption CTA. While Assumption SE allows $\Delta_{k}Y_{it}(d)$ to depend on $(W_{i,t+k},W_{it},\overline{W}_{i})$, a TWFE regression predicts outcome changes using only $\Delta_{k}W_{it}$. Assumption S eliminates potential omitted variable bias from this discrepancy. Together, Assumptions SE and S imply Assumption CTA.
Theorem B3.
Under Assumptions PO, V, LH, LC, SE, and S,
$\beta_{\text{FE}}^{D}=\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(x)\psi_{it}^{*}(x)\right]dx,$
where the weight function is given by:
$\psi_{it}^{*}(d)\equiv\frac{\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}_{D_{is}\geq d}R_{is}+\frac{1}{T}\sum_{s=1}^{T}E\left[\mathbbm{1}_{D_{it}\geq d}R_{it}-\mathbbm{1}_{D_{is}\geq d}R_{is}|W_{it}-W_{is},C_{i}\right]}{\sum_{s=1}^{T}E\left[D_{is}R_{is}\right]}$ (B.1)
and satisfies $\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[\psi_{it}^{*}(x)\right]dx=1$. In addition,
$\sum_{t=1}^{T}E\left[\psi_{it}^{*}(d)^{2}\right]\leq\sum_{t=1}^{T}E\left[\left(\psi_{it}^{*}(d)+g(d)R_{it}\right)^{2}\right]\thinspace\thinspace\thinspace\text{for all }\thinspace\text{d$\in$($\underline{d}$,$\overline{d}$)}$ (B.2)
holds for any $g:(\underline{d},\overline{d})\mapsto\mathbb{R}$. $\psi_{it}^{*}(d)$ can also be expressed as
$\psi_{it}^{*}(d)=\frac{1}{T}\sum_{s=1}^{T}\psi_{is}(d;d^{0})+\frac{1}{T}\sum_{s=1}^{T}E\left[\psi_{it}(d;d^{0})-\psi_{is}(d;d^{0})|W_{it}-W_{is},C_{i}\right],$ (B.3)
where $\psi_{it}(d;d^{0})$ is the weight function defined in Theorem B2 for any $d^{0}\in(\underline{d},\overline{d})$.
Theorem B3 expresses the coefficient $\beta_{\text{FE}}^{D}$ as a weighted average of treatment effects $Y_{it}^{\prime}(d)$, with weights summing to one but still possibly negative, much like Theorem B2. However, its weight function $\psi_{it}^{*}(d)$ does not depend on an arbitrary baseline level $d^{*}$. While any function proportional to $R_{it}$ can still be added to $\psi_{it}^{*}(d)$ without affecting the weighted average, doing so always increases the weight’s square norm as shown in (B.2).
As observed in (B.1), the weight function $\psi_{it}^{*}(d)$ is additive in two components, where the first component is constant within the same unit $i$ and the second component is constant within the same period $t$ conditional on covariates. Moreover, as expressed in (B.3), $\psi_{it}^{*}(d)$ can be obtained by collapsing any possible weight function $\psi_{it}(d;d^{0})$ into the unit-specific and the period-specific terms in exactly the same manner.
Considering the simplest case with no covariates is helpful for offering an intuition for the weight function. In this case, the weight function is given by
$\psi_{it}^{*}(d)=\frac{\frac{1}{T}\sum_{s=1}^{T}\left(\mathbbm{1}_{D_{is}\geq d}R_{is}-E\left[\mathbbm{1}_{D_{is}\geq d}R_{is}\right]\right)+E\left[\mathbbm{1}_{D_{it}\geq d}R_{it}\right]}{\sum_{s=1}^{T}E\left[D_{is}R_{is}\right]},$
which is a predicted value from a regression of $\psi_{it}(d;d^{0})$ on additive unit and period fixed effects. Given that a common trends assumption relies on treatment effects $Y_{it}^{\prime}(d)$ being additive in unit and period effects, any weight variation orthogonal to unit and period effects is orthogonal to $Y_{it}^{\prime}(d)$. Projecting the weight function onto unit and period effects eliminates this redundant weight variation. [^12] See Ishimaru (2024, Section III.A) for a discussion in a related context, in which an instrumental variables regression employs a DID-type identification strategy.
While the set of assumptions in Theorem 4 also holds in Theorem B3, using the weight function $\omega_{it}$ in Theorem 4 can be misleading for defining the weight on each observation. In particular, the total weight on an observation ($i,t$) aggregated over $d\in(\underline{d},\overline{d})$ in Theorem B3 is given by:
$\omega_{it}^{*}\equiv\int_{\underline{d}}^{\overline{d}}\psi_{it}^{*}(x)dx=\frac{\frac{1}{T}\sum_{s=1}^{T}D_{is}R_{is}+\frac{1}{T}\sum_{s=1}^{T}E\left[D_{it}R_{it}-D_{is}R_{is}|W_{it}-W_{is},C_{i}\right]}{\sum_{s=1}^{T}E\left[D_{is}R_{is}\right]}\\ =\frac{1}{T}\sum_{s=1}^{T}\omega_{is}+\frac{1}{T}\sum_{s=1}^{T}E\left[\omega_{it}-\omega_{is}|W_{it}-W_{is},C_{i}\right].$ (B.4)
This weight projection eliminates redundant variation in $\omega_{it}$ exactly in the same manner as in (B.3). This projection holds for any baseline level $d^{0}$ chosen for defining $\omega_{it}$. For example, in a binary treatment setting under Assumption CTA, it is well known that selecting $d^{0}=0$ gives treated observations nonzero (potentially negative) weights, while $d^{0}=1$ gives untreated observations such weights. Applying the projection in equation (B.4) to either weighting scheme yields $\omega_{it}^{*}$.
B.3 Proofs
Proof of Theorem B1
Under Assumptions PO–D and CE–D, observed outcome changes decompose as:
$\Delta_{k}Y_{it}=\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})+\sum_{\ell=0}^{L}\gamma_{\ell}\Delta_{k}D_{i,t-\ell}.$
Then, the numerator in (11) is given by
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ +\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{i,t-\ell}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]\\ =\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ +\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{i,t-\ell}\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ =\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Cov\left(\Delta_{k}D_{i,t-\ell},\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right],$
where the second equality uses Assumption LH and the third equality uses Assumption CT–D. In addition, the denominator in (11) is given by
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]=\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ =\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right],$
where the first equality uses Assumption LH.
The numerator in (A.11) is given by
$\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]+\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{i,t-\ell}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FD},k}^{W}\right)\right]\\ =\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d^{0},\ldots,d^{0})\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ +\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{i,t-\ell}\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ =\sum_{\ell=0}^{L}\gamma_{\ell}\sum_{t=1}^{T-k}E\left[Cov\left(\Delta_{k}D_{i,t-\ell},\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right],$
where the second equality uses Assumption LH and the third equality uses Assumption CT–D. In addition, the denominator in (A.11) is given by
$\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}\right)\right]=\sum_{t=1}^{T-k}E\left[\Delta_{k}D_{it}\left(\Delta_{k}D_{it}-E[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}]\right)\right]\\ =\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right)\right],$
where the first equality uses Assumption LH.
Proof of Theorem B2
Given the definition of $R_{it}$ and Assumption LH,
$\Delta_{k}R_{it}$ $=\Delta_{k}\widetilde{D}_{it}-\Delta_{k}\widetilde{W}_{it}^{\prime}\delta_{\text{FE}}^{W}$
$=\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right].$ (B.5)
For any $d\in(\underline{d},\overline{d})$,
$\sum_{t=1}^{T}E\left[Y_{it}(d)R_{it}\right]$ $=$ $\frac{1}{T}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d)\Delta_{k}R_{it}\right]$ (B.6)
$=$ $\frac{1}{T}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}(d)\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i}\right]\right)\right]$
$=$ $0.$
The first equality uses (A.1) and $\overline{R}_{i}=0$. The second equality uses (B.5). The third equality follows from Assumption CTA. Under Assumption LC, (B.6) also implies
$\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(d)R_{it}\right]=0$ (B.7)
for any $d\in(\underline{d},\overline{d})$.
The numerator of (11) is given by
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\Delta_{k}Y_{it}\Delta_{k}R_{it}\right]$ $=$ $T\sum_{t=1}^{T}E\left[Y_{it}R_{it}\right]$ (B.8)
$=$ $T\sum_{t=1}^{T}E\left[\left(Y_{it}(D_{it})-Y_{it}(d^{0})\right)R_{it}\right]$
$=$ $T\sum_{t=1}^{T}E\left[\int_{\underline{d}}^{\overline{d}}Y_{it}^{\prime}(x)\left(\mathbbm{1}_{D_{it}\geq x}-\mathbbm{1}_{d^{0}\geq x}\right)dxR_{it}\right]$
$=$ $T\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(x)\left(\mathbbm{1}_{D_{it}\geq x}-\mathbbm{1}_{d^{0}\geq x}\right)R_{it}\right]dx.$
The first equality uses (A.1) and $\overline{R}_{i}=0$, and the second equality uses (B.6) and Assumption PO. The third equality follows from the fundamental theorem of calculus and an identity that $\mathbbm{1}_{D_{it}\geq x}-\mathbbm{1}_{d^{0}\geq x}$ is equal to $1$ if $d^{0}<x\leq D_{it}$, to 1-1 if $D_{it}<x\leq d^{0}$, and to 0 otherwise. The fourth equality follows from Fubini’s theorem. Note that the absolute integrability condition required for using the theorem is satisfied because Assumption LC and the Cauchy–Schwarz inequality imply
$E\left[\int_{\underline{d}}^{\overline{d}}\left|Y_{it}^{\prime}(x)\left(\mathbbm{1}_{D_{it}\geq x}-\mathbbm{1}_{d^{0}\geq x}\right)R_{it}\right|dx\right]$ $\leq$ $E\left[M_{it}\cdot\left|D_{it}-d^{0}\right|\cdot\left|R_{it}\right|\right]$
$\leq$ $\sqrt{E\left[M_{it}^{2}\left(D_{it}-d^{0}\right)^{2}\right]E\left[R_{it}^{2}\right]}$
$<$ $\infty.$
Combining (B.7) and (B.8) yields the weighted average expression in Theorem B2.
In addition,
$\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[\psi_{it}(x;d^{0})\right]dx$ $=\sum_{t=1}^{T}E\left[\int_{\underline{d}}^{\overline{d}}\psi_{it}(x;d^{0})dx\right]$
$=\frac{\sum_{t=1}^{T}E\left[\left(D_{it}-d^{0}\right)R_{it}\right]}{\sum_{t=1}^{T}E\left[D_{it}R_{it}\right]}$
$=1,$
where the first equality uses Fubini’s theorem and the last equality follows from $E[R_{it}]=0$.
Proof of Theorem B3
Since the set of assumptions in Theorem B2 is still satisfied,
$\beta_{FE}^{D}=\int_{\underline{d}}^{\overline{d}}\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(x)\psi_{it}(x;d^{0})\right]dx.$
For any $d\in(\underline{d},\overline{d})$,
$\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(d)\left(\psi_{it}(d;d^{0})-\frac{1}{T}\sum_{s=1}^{T}\psi_{is}(d;d^{0})\right)\right]$
$=$ $\frac{1}{T}\sum_{s>t}E\left[\left(Y_{is}^{\prime}(d)-Y_{it}^{\prime}(d)\right)\left(\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})\right)\right]$
$=$ $\frac{1}{T}\sum_{s>t}E\left[E\left[Y_{is}^{\prime}(d)-Y_{it}^{\prime}(d)|D_{is},D_{it},\overline{D}_{i},W_{is},W_{it},\overline{W}_{i},C_{i}\right]\left(\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})\right)\right]$
$=$ $\frac{1}{T}\sum_{s>t}E\left[E\left[Y_{is}^{\prime}(d)-Y_{it}^{\prime}(d)|W_{is}-W_{it},C_{i}\right]\left(\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})\right)\right]$
$=$ $\frac{1}{T}\sum_{s>t}E\left[\left(Y_{is}^{\prime}(d)-Y_{it}^{\prime}(d)\right)E\left[\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})|W_{is}-W_{it},C_{i}\right]\right].$ (B.9)
The first equality uses (A.1). The second equality follows from the law of iterated expectations, since $\psi_{it}(d;d^{0})$ is determined by $(D_{it},\overline{D}_{i},W_{it},\overline{W}_{i},C_{i})$. The third equality uses Assumptions SE and S. The last equality follows from the law of iterated expectations.
Letting $A(s,t)=E\left[\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})|W_{is}-W_{it},C_{i}\right]$ and $b_{t}=Y_{it}(d)$ in (B.9) yields
$\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(d)\psi_{it}(d;d^{0})\right]=\sum_{t=1}^{T}E\left[Y_{it}^{\prime}(d)\frac{1}{T}\sum_{s=1}^{T}\left(\psi_{is}(d;d^{0})+E\left[\psi_{it}(d;d^{0})-\psi_{is}(d;d^{0})|W_{it}-W_{is},C_{i}\right]\right)\right].$
This proves (B.3).
In addition,
$\frac{1}{T}\sum_{s=1}^{T}\psi_{is}(d;d^{0})=\frac{\frac{1}{T}\sum_{s=1}^{T}\left(\mathbbm{1}_{D_{is}\geq d}-\mathbbm{1}_{d^{0}\geq d}\right)R_{is}}{\sum_{t^{\prime}=1}^{T}E\left[D_{it^{\prime}}R_{it}\right]}=\frac{\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}_{D_{is}\geq d}R_{is}}{\sum_{t^{\prime}=1}^{T}E\left[D_{it^{\prime}}R_{it}\right]}$
follows from $\frac{1}{T}\sum_{s=1}^{T}R_{is}=0$ and
$E\left[\psi_{it}(d;d^{0})-\psi_{is}(d;d^{0})|W_{it}-W_{is},C_{i}\right]$
$=$ $\frac{E\left[\mathbbm{1}_{D_{it}\geq d}R_{it}-\mathbbm{1}_{D_{is}\geq d}R_{is}-(R_{it}-R_{is})\mathbbm{1}_{d^{0}\geq d}|W_{it}-W_{is},C_{i}\right]}{\sum_{t^{\prime}=1}^{T}E\left[D_{it^{\prime}}R_{it}\right]}$
$=$ $\frac{E\left[\mathbbm{1}_{D_{is}\geq d}R_{is}-\mathbbm{1}_{D_{it}\geq d}R_{it}|W_{it}-W_{is},C_{i}\right]}{\sum_{t^{\prime}=1}^{T}E\left[D_{it^{\prime}}R_{it}\right]}$
follows from (B.5). This proves (B.1).
Finally,
$\sum_{t=1}^{T}E\left[\psi_{it}^{*}(d)R_{it}\right]$ $=\frac{1}{T}\sum_{s=1}^{T}\sum_{t=1}^{T}E\left[R_{it}\left(\psi_{is}(d;d^{0})+E\left[\psi_{it}(d;d^{0})-\psi_{is}(d;d^{0})|W_{it}-W_{is},C_{i}\right]\right)\right]$
$=\frac{1}{T}\sum_{s=1}^{T}\sum_{t=1}^{T}E\left[R_{it}E\left[\psi_{it}(d;d^{0})-\psi_{is}(d;d^{0})|W_{it}-W_{is},C_{i}\right]\right]$
$=\frac{1}{T}\sum_{s>t}E\left[\left(R_{is}-R_{it}\right)E\left[\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})|W_{is}-W_{it},C_{i}\right]\right]$
$=\frac{1}{T}\sum_{s>t}E\left[E\left[R_{is}-R_{it}|W_{is}-W_{it},C_{i}\right]\left(\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})\right)\right]$
$=0,$
where the first equality uses (B.3), the second equality follows from $\sum_{t=1}^{T}R_{it}=0$, the third equality uses (A.2) by letting $A(s,t)=E\left[\psi_{is}(d;d^{0})-\psi_{it}(d;d^{0})|W_{is}-W_{it},C_{i}\right]$ and $b_{t}=R_{it}$, the fourth equality uses the law of iterated expectations, and the last equality uses Assumption (B.5). This proves (B.2).
Appendix C Results in Unbalanced Panels
The main paper assumes a balanced panel setting, in which each unit–period combination has exactly one observation. This section discusses how the results in the main paper extends to or differs in a more general setting. Proofs of additional theoretical results are provided at the end of this section.
C.1 Equivalence of Least-Square Objectives
Without assuming a balanced panel, the least-squares objective in (4) should be rewritten as
$\underset{\beta,\{\alpha_{i}\}_{i=1}^{N},\{\gamma_{t},\mu_{t}\}_{t=1}^{T}}{\min}\sum_{i=1}^{N}\sum_{t=1}^{T}B_{it}\left(Y_{it}-X_{it}^{\prime}\beta-\alpha_{i}-C_{i}^{\prime}\gamma_{t}-\mu_{t}\right)^{2}.$ (C.1)
In an unbalanced panel, $B_{it}$ is an observability indicator hat takes a value of 1 if an observation $(i,t)$ is available and a value of 0 if the observation is missing. More generally, $B_{it}$ can be any real number. For example, a TWFE regression may use repeated cross-section data with $B_{it}$ observations in group $i$ at period $t$. The equivalence between TWFE and pooled FD regressions naturally extends to these cases.
Theorem C1.
The coefficient on $X_{it}$ given by (C.1) is algebraically identical to the coefficient on $\Delta_{k}X_{it}$ given by the following least-squares problem.
$\underset{\beta,\left\{\gamma_{t},\mu_{t}\right\}_{t=1}^{T}}{\min}$ $\sum_{k=1}^{T-1}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\Delta_{k}\gamma_{t}-\Delta_{k}\mu_{t}\right)^{2}.$ (C.2)
C.2 Weighted-Average Relationship
The least-square problem (C.2) is distinct from
$\underset{\beta,\left\{\gamma_{k,t}^{*}\right\}_{k,t}}{\min}\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{k,t}^{*}-\mu_{k,t}^{*}\right)^{2}.$ (C.3)
While (C.2) includes fixed effects for starting and ending periods in an additive manner, (C.3) accounts for fixed effects associated with all possible combinations of starting and ending periods. These separability restrictions, $\gamma_{k,t}^{*}=\gamma_{t+k}-\gamma_{t}$ and $\mu_{k,t}^{*}=\mu_{t+k}-\mu_{t}$ , do not bind in a balanced panel, as seen in the proof of Theorem 2. However, this property does not extend to unbalanced panels.
As a result, even when $X_{it}$ is univariate, unlike in Theorem 2, the TWFE coefficient $\widehat{\beta}_{\text{FE}}$ given by (C.1) or (C.2) is not exactly identical to a convex combination of FD coefficients $\widehat{\beta}_{\text{FD},k}$ that are given by solving
$\underset{\beta,\left\{\gamma_{t}^{*},\mu_{t}^{*}\right\}_{t=1}^{T-k}}{\min}\sum_{i=1}^{N}\sum_{t=1}^{T-k}\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{t}^{*}-\mu_{t}^{*}\right)^{2}$ (C.4)
for $k=1,\ldots,T-1$. One possible way to restore a weighted-average relationship is to use (C.3) instead of (C.1) or (C.2) for producing the TWFE coefficient.
C.3 Causal Interpretation
Now, I consider causal interpretation of the population TWFE coefficient that arises from the regression (C.1). For simplicity, I focus on an unbalanced panel in which $B_{it}$ is either 0 or $1$ and $\overline{B}_{i}>0$. Furthermore, I focus on the setting with a baseline treatment level, which corresponds to the one in Section 3.1. The required assumptions for causal interpretation of the TWFE coefficient should be modified as follows.
Assumption CT-u.
(Conditional Common Trends) There exists $d^{0}\in(\underline{d},\overline{d})$ such that
$E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}D_{it},\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]=E\left[\Delta_{k}Y_{it}(d^{0})|\Delta_{k}W_{it},C_{i},B_{it}=B_{i,t+k}=1\right]$
for any $(k,t)$ with $1\leq t<t+k\leq T$.
Assumption V-u.
(Variation) $\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[Var\left(\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right)\right]>0$.
Assumption LH-u.
(Linearity and Time Homogeneity) For any $(k,t)$ with $1\leq t<t+k\leq T$,
$E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]=\Delta_{k}W_{it}^{\prime}\delta^{W}+C_{i}^{\prime}\left(\delta_{t+k}^{C}-\delta_{t}^{C}\right)+\left(\delta_{t+k}^{I}-\delta_{t}^{I}\right).$
where the coefficient $\delta^{W}$ does not vary across $(k,t)$.
These assumptions differ from the corresponding assumptions (CT, V, and LH) in Section 3 in three ways. First, the assumptions about the potential outcome change $\Delta_{k}Y_{it}(d^{0})$ and the treatment change $\Delta_{k}D_{it}$ are now conditional on the changes being observed, $B_{it}=B_{i,t+k}=1$. Second, the observation rate $\overline{B}_{i}$ has to be exogenous, although this requirement can be removed when time-invariant covariates $C_{i}$ include $\overline{B}_{i}$. Third, the coefficient on $C_{i}$ and the intercept coefficient in Assumption LH-u are less flexible than Assumption LH, in which the coefficients can depend on $(t,k)$ in an arbitrary manner.
Theorem C2.
Under Assumptions PO, CT-u, V-u, and LH-u,
$\beta_{\text{FE}}^{D}=\sum_{t=1}^{T}E\left[\tau_{it}\omega_{it}\right],$
where $\tau_{it}$ is a per-unit effect defined in Theorem 4. The weights satisfy $\sum_{t=1}^{T}E\left[\omega_{it}\right]=1$ and
$\omega_{it}\propto\frac{B_{it}}{\overline{B}_{i}}\left(D_{it}-d^{0}\right)\sum_{s=1}^{T}B_{is}\left(D_{it}-D_{is}-E\left[D_{it}-D_{is}|W_{it}-W_{is},C_{i},\overline{B}_{i},B_{it}=B_{is}=1\right]\right).$ (C.5)
This theorem, much like Theorem 4, expresses the coefficient $\beta_{\text{FE}}^{D}$ as a weighted average of per-unit treatment effects $\tau_{it}$, with weights summing to one but possibly negative. The weight function is similar to Theorem 4 but is adjusted for observability $B_{it}$.
C.4 Proofs
Proof of Theorem C1
Suppose a sequence $\{w_{t}\}_{t=1}^{T}$ satisfies $\sum_{t=1}^{T}w_{t}=T$ and define $\overline{a}^{w}\equiv\frac{1}{T}\sum_{t=1}^{T}w_{t}a_{t}$ for another sequence $\{a_{t}\}_{t=1}^{T}$. Basic algebra yields
$\frac{1}{T}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}w_{t+k}w_{t}\left(\Delta_{k}a_{t}\right)^{2}=\frac{1}{2T}\sum_{s=1}^{T}\sum_{t=1}^{T}w_{s}w_{t}\left(a_{s}-a_{t}\right)^{2}=\frac{1}{2T}\sum_{s=1}^{T}\sum_{t=1}^{T}w_{s}w_{t}\left(a_{s}-\overline{a}^{w}+\overline{a}^{w}-a_{t}\right)^{2}\\ =\frac{1}{2T}\sum_{s=1}^{T}\sum_{t=1}^{T}w_{s}w_{t}\left\{\left(a_{s}-\overline{a}^{w}\right)^{2}+\left(a_{t}-\overline{a}^{w}\right)^{2}-2\left(a_{s}-\overline{a}^{w}\right)\left(a_{t}-\overline{a}^{w}\right)\right\}=\sum_{t=1}^{T}w_{t}\left(a_{t}-\overline{a}^{w}\right)^{2}\\ =\sum_{t=1}^{T}w_{t}a_{t}^{2}-\left(\overline{a}^{w}\right)^{2}.$ (C.6)
Exploiting the algebraic identity (C.6) by letting $w_{t}=\frac{B_{it}}{\overline{B}_{i}}$ and $a_{t}=Y_{it}-X_{it}^{\prime}\beta-\alpha_{i}-C_{i}^{\prime}\gamma_{t}-\mu_{t}$ for each $i=1,\ldots,N$, the objective function (C.1) can be expressed as
$\sum_{i=1}^{N}\sum_{t=1}^{T}B_{it}\left(Y_{it}-X_{it}^{\prime}\beta-\alpha_{i}-C_{i}^{\prime}\gamma_{t}-\mu_{t}\right)^{2}\\ =\frac{1}{T}\sum_{i=1}^{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\frac{B_{i,t+k}B_{it}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\Delta_{k}X_{it}^{\prime}\beta-C_{i}^{\prime}\Delta_{k}\gamma_{t}-\Delta_{k}\mu_{t}\right)^{2}\\ +\sum_{i=1}^{N}\overline{B}_{i}\left(\sum_{t=1}^{T}\frac{B_{it}}{T\overline{B}_{i}}\left(Y_{it}-X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{t}-\mu_{t}\right)-\alpha_{i}\right)^{2}.$ (C.7)
Since choosing $\beta$ and $\{\gamma_{t},\mu_{t}\}_{t=1}^{T}$ that solve (C.2) and letting $\alpha_{i}=\sum_{t=1}^{T}\frac{B_{it}}{T\overline{B}_{i}}\left(Y_{it}-X_{it}^{\prime}\beta-C_{i}^{\prime}\gamma_{t}-\mu_{t}\right)$ for each $i=1,\ldots,N$ results in minimizing (C.7), the two objectives (C.1) and (C.2) yields the same estimates of $\beta$.
Proof of Theorem C2
Suppose $X_{it}$ consists of a scalar treatment variable $D_{it}$ and a vector of time-varying covariates $W_{it}$ in (C.1). Due to the equivalence between (C.1) and (C.2), the population TWFE coefficient on $D_{it}$ is given by solving
$\underset{\beta^{D},\beta^{W},\left\{\gamma_{t},\mu_{t}\right\}_{t=1}^{T}}{\min}$ $\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\beta^{D}\Delta_{k}D_{it}-\Delta_{k}W_{it}^{\prime}\beta^{W}-C_{i}^{\prime}\Delta_{k}\gamma_{t}-\Delta_{k}\mu_{t}\right)^{2}\right].$ (C.8)
Assumption LH-u and the FWL theorem imply that the same $\beta^{D}$ as (C.8) can be recovered by solving
$\underset{\beta^{D}}{\min}$ $\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left\{\Delta_{k}Y_{it}-\beta^{D}\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right\}^{2}\right].$ (C.9)
This regression yields
$\beta^{D}=\frac{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\Delta_{k}Y_{it}\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right]}{\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\Delta_{k}D_{it}\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right]}.$ (C.10)
The numerator in (C.10) can be expressed as
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\Delta_{k}Y_{it}\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right]$
$=$ $\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\left(\Delta_{k}Y_{it}-\Delta_{k}Y_{it}(d^{0})\right)\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right]$
$=$ $\sum_{s=1}^{T}\sum_{t=1}^{T}E\left[\frac{B_{it}B_{is}}{\overline{B}_{i}}\left(Y_{it}(D_{it})-Y_{it}(d^{0})\right)\left(D_{is}-D_{it}-E\left[D_{is}-D_{it}|W_{is}-W_{it},C_{i},\overline{B}_{i},B_{it}=B_{is}=1\right]\right)\right]$
$=$ $\sum_{t=1}^{T}E\left[\tau_{it}\frac{B_{it}}{\overline{B}_{i}}\left(D_{it}-d^{0}\right)\sum_{s=1}^{T}B_{is}\left(D_{it}-D_{is}-E\left[D_{it}-D_{is}|W_{it}-W_{is},C_{i},\overline{B}_{i},B_{it}=B_{is}=1\right]\right)\right].$ (C.11)
Note that the first equality uses Assumption CT-u, the second equality follows from Assumption PO and (A.2), and the last equality uses the definition of $\tau_{it}$.
The denominator in (C.10) is strictly positive due to Assumption V-u. Also, using the transformation applied to the numerator by letting $Y_{it}(d)=d$, the denominator in (C.10) can be expressed as
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}E\left[\frac{B_{it}B_{i,t+k}}{\overline{B}_{i}}\Delta_{k}D_{it}\left(\Delta_{k}D_{it}-E\left[\Delta_{k}D_{it}|\Delta_{k}W_{it},C_{i},\overline{B}_{i},B_{it}=B_{i,t+k}=1\right]\right)\right]$
$=$ $\sum_{t=1}^{T}E\left[\frac{B_{it}}{\overline{B}_{i}}\left(D_{it}-d^{0}\right)\sum_{s=1}^{T}B_{is}\left(D_{it}-D_{is}-E\left[D_{it}-D_{is}|W_{it}-W_{is},C_{i},\overline{B}_{i},B_{it}=B_{is}=1\right]\right)\right].$ (C.12)
Combining (C.11) and (C.12) yields the weight expression (C.5) and $\sum_{t=1}^{T}E\left[\omega_{it}\right]=1$.
Appendix D Connection of Theorem 1 with Prior Work
This section explores how Theorem 1, which links TWFE and pooled FD regressions via a U-statistics identity, connects to prior DID decomposition results, notably Goodman-Bacon (2021) and Li and Strezhnev (2024). It extends the period-based expansion in Theorem 1 to cross-unit differences and situates earlier findings within a broader numerical framework.
Cross-Unit Expansion of TWFE
Theorem 1 derives TWFE coefficients from a least-squares problem pooling $k$-period differences over time. Given TWFE’s symmetry across units and periods, this approach extends to differences across units, as shown below.
Theorem D1.
Let $\widehat{\beta}_{\text{FE}}$ be the coefficient on $X_{it}$ from the TWFE least-squares problem defined in (4). Then $\beta=\widehat{\beta}_{\text{FE}}$ is also solves:
$\underset{\beta,\{\gamma_{t}\}_{t=1}^{T}}{\min}$ $\sum_{i<j}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left\{\left(\Delta_{k}Y_{jt}-\Delta_{k}Y_{it}\right)-\left(\Delta_{k}X_{jt}-\Delta_{k}X_{it}\right)^{\prime}\beta-\left(C_{j}-C_{i}\right)^{\prime}\Delta_{k}\gamma_{t}\right\}^{2}.$ (D.1)
Proof.
This result builds on the proof of Theorem 1 presented in Appendix A. Applying the U-statistics identity (A.1) to the cross-sectional summation across $i=1,\ldots,N$ in (A.3) yields:
$\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\sum_{i=1}^{N}\left(\Delta_{k}e_{it}\right)^{2}=\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left\{\sum_{i=1}^{N}\left(\Delta_{k}e_{it}-\frac{1}{N}\sum_{j=1}^{N}\Delta_{k}e_{jt}\right)^{2}+N\left(\frac{1}{N}\sum_{i=1}^{N}\Delta_{k}e_{it}\right)^{2}\right\}\\ =\frac{1}{N}\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\sum_{1\leq i<j\leq N}\left(\Delta_{k}e_{jt}-\Delta_{k}e_{it}\right)^{2}+N\sum_{k=1}^{T-1}\sum_{t=1}^{T-k}\left(\frac{1}{N}\sum_{i=1}^{N}\Delta_{k}e_{it}\right)^{2}.$ (D.2)
[…]
The least-squares problem (D.1) can be interpreted as a regression of a DID in $Y_{it}$ across units and periods on the corresponding DID in $X_{it}$, controlling for a difference in time-invariant covariates $C_{i}$. For a scalar $X_{it}$ without covariates, Theorem D1 implies:
$\widehat{\beta}_{\text{FE}}=$ $\frac{\sum_{s>t,j>i}\left\{(Y_{js}-Y_{jt})-(Y_{is}-Y_{it})\right\}\left\{(X_{js}-X_{jt})-(X_{is}-X_{it})\right\}}{\sum_{s>t,j>i}\left\{(X_{js}-X_{jt})-(X_{is}-X_{it})\right\}^{2}}.$ (D.3)
In a staggered adoption design with binary $X_{it}$, each nonzero term in the numerator corresponds to a two-period, two-unit DID: it equals $(Y_{is}-Y_{it})-(Y_{js}-Y_{jt})$ if only unit $i$ starts treatment between periods $t$ and $s$, or the reserve if only $j$ does. Thus, $\widehat{\beta}_{\text{FE}}$ averages all such DID comparisons.
Relation to Prior Work
This representation aligns with Goodman-Bacon (2021) and Li and Strezhnev (2024), who decompose TWFE estimators in binary treatment settings into averages of two-unit, two-period DID comparisons. Their approaches rely on algebraic methods tailored to binary treatment settings. While Goodman-Bacon (2021) employs a similar identity as a lemma, he uses it for an expansion across units but not periods, and does not present it as a U-statistics result. By contrast, my approach explicitly leverages the U-statistics identity to uncover a general structure that applies to binary, discrete, or continuous regressors, with or without covariates, revealing their results as special cases.
To further illustrate the link, I demonstrate that the weighted-average expression in Goodman-Bacon (2021) results from aggregating (D.3) by treatment timing. Define the treatment as $X_{it}=\mathbbm{1}_{t>g_{i}}$, where $g_{i}\in\{1,\ldots,T\}$ represents the last period in which the unit $i$ is untreated, with $g_{i}=T$ for the never-treated group (denoted $U$ in Goodman-Bacon, 2021). Without loss of generality, suppose that units $i=1,\ldots,N$ are ordered so that $g_{j}\leq g_{i}$ for any $j>i$. Let $n_{g}$ be the number of units that belongs to a group $g$, where $\sum_{g=1}^{T}n_{g}=N$. Define
$\overline{Y}_{gt}=\frac{1}{n_{g}}\sum_{i:g_{i}=g}Y_{it}$
to be the group mean of the outcome. Given these,
$\sum_{s>t,j>i}\left\{(Y_{js}-Y_{jt})-(Y_{is}-Y_{it})\right\}\left\{(X_{js}-X_{jt})-(X_{is}-X_{it})\right\}$
$=$ $\sum_{s>t,h>g}n_{g}n_{h}\left\{(\overline{Y}_{gs}-\overline{Y}_{gt})-(\overline{Y}_{hs}-\overline{Y}_{ht})\right\}\left(\mathbbm{1}_{t\leq g<s}-\mathbbm{1}_{t\leq h<s}\right)$
$=$ $\frac{1}{N}\sum_{t\leq g<s\leq h}n_{g}n_{h}\left\{(\overline{Y}_{gs}-\overline{Y}_{gt})-(\overline{Y}_{hs}-\overline{Y}_{ht})\right\}-\frac{1}{N}\sum_{g<t\leq h<s}n_{g}n_{h}\left\{(\overline{Y}_{gs}-\overline{Y}_{gt})-(\overline{Y}_{hs}-\overline{Y}_{ht})\right\},$
where the last equality follows from the identity that $\mathbbm{1}_{t\leq g<s}-\mathbbm{1}_{t\leq h<s}$ is equal to $1$ if $t\leq g<s\leq h$, to 1-1 if $g<t\leq h<s$, and to 0 otherwise.
Then, following the notations in Goodman-Bacon (2021), for any two groups $g$ and $h$, let
$\overline{Y}_{g}^{POST(h)}=\frac{1}{T-h}\sum_{t=h+1}^{T}\overline{Y}_{gt}$
be the mean outcome of the group $g$ during periods in which the group $h<T$ is treated, and let
$\overline{Y}_{g}^{PRE(h)}=\frac{1}{h}\sum_{t=1}^{h}\overline{Y}_{gt}$
be the mean outcome of the group $g$ during periods in which the group $h$ is untreated. In addition, for $g<h$, let
$\overline{Y}_{m}^{MID(g,h)}=\frac{1}{h-g}\sum_{t=g+1}^{h}\overline{Y}_{mt}$
be the mean outcome of a group $m\in\{g,h\}$ during periods in which the group $g$ is treated and the group $h$ is untreated.
Using these definitions,
$\sum_{s>t,j>i}\left\{(Y_{js}-Y_{jt})-(Y_{is}-Y_{it})\right\}\left\{(X_{js}-X_{jt})-(X_{is}-X_{it})\right\}$
$=$ $\frac{1}{N}\sum_{g=1}^{T-1}\sum_{h=g+1}^{T}\sum_{t=1}^{g}\sum_{s=g+1}^{h}n_{g}n_{h}\left\{(\overline{Y}_{gs}-\overline{Y}_{gt})-(\overline{Y}_{hs}-\overline{Y}_{ht})\right\}$
$-\frac{1}{N}\sum_{g=1}^{T-2}\sum_{h=g+1}^{T-1}\sum_{t=g+1}^{h}\sum_{s=h+1}^{T}n_{g}n_{h}\left\{(\overline{Y}_{gs}-\overline{Y}_{gt})-(\overline{Y}_{hs}-\overline{Y}_{ht})\right\}$
$=$ $\frac{1}{N}\sum_{g=1}^{T-1}\sum_{h=g+1}^{T}g(h-g)n_{g}n_{h}\left\{(\overline{Y}_{g}^{MID(g,h)}-\overline{Y}_{g}^{PRE(g)})-(\overline{Y}_{h}^{MID(g,h)}-\overline{Y}_{h}^{PRE(g)})\right\}$
$+\frac{1}{N}\sum_{g=1}^{T-2}\sum_{h=g+1}^{T-1}(h-g)(T-h)n_{g}n_{h}\left\{(\overline{Y}_{h}^{POST(h)}-\overline{Y}_{h}^{MID(g,h)})-(\overline{Y}_{g}^{POST(h)}-\overline{Y}_{g}^{MID(g,h)})\right\}.$ (D.4)
Note that $(\overline{Y}_{g}^{MID(g,h)}-\overline{Y}_{g}^{PRE(g)})-(\overline{Y}_{h}^{MID(g,h)}-\overline{Y}_{h}^{PRE(g)})$ is a DID comparison between an early-treated group $g$ and a late-treated (or never-treated) group $h$ across the periods in which only $g$ is treated and the periods in which both are untreated, the not-yet-treated $h$ serving as the control group. In addition, $(\overline{Y}_{h}^{POST(h)}-\overline{Y}_{h}^{MID(g,h)})-(\overline{Y}_{g}^{POST(h)}-\overline{Y}_{g}^{MID(g,h)})$ is a DID comparison between a late-treated group $h$ and an early-treated group $g$ across the periods in which both $g$ and $h$ are treated and the period in which only $g$ is treated, the already-treated $g$ serving as the control group.
Similarly, the denominator in (D.3) satisfies
$\sum_{s>t,j>i}\left\{(X_{js}-X_{jt})-(X_{is}-X_{it})\right\}^{2}=\frac{1}{N}\sum_{g=1}^{T-1}\sum_{h=g+1}^{T}g(h-g)n_{g}n_{h}+\frac{1}{N}\sum_{g=1}^{T-2}\sum_{h=g+1}^{T-1}(h-g)(T-h)n_{g}n_{h}.$ (D.5)
Combining (D.4) and (D.5) yields
$\widehat{\beta}^{FE}=$ $\sum_{g=1}^{T-1}w_{gT}\left\{(\overline{Y}_{g}^{POST(g)}-\overline{Y}_{g}^{PRE(g)})-(\overline{Y}_{T}^{POST(g)}-\overline{Y}_{T}^{PRE(g)})\right\}$
$+$ $\sum_{g=1}^{T-2}\sum_{h=g+1}^{T-1}w_{gh}\left[\frac{g}{T-h+g}\left\{(\overline{Y}_{g}^{MID(g,h)}-\overline{Y}_{g}^{PRE(g)})-(\overline{Y}_{h}^{MID(g,h)}-\overline{Y}_{h}^{PRE(g)})\right\}\right$.
$\left.+\frac{T-h}{T-h+g}\left\{(\overline{Y}_{h}^{POST(h)}-\overline{Y}_{h}^{MID(g,h)})-(\overline{Y}_{g}^{POST(h)}-\overline{Y}_{g}^{MID(g,h)})\right\}\right],$ (D.6)
where the weights are defined as
$w_{gh}=\frac{n_{g}n_{h}(h-g)(T-h+g)}{\sum_{g^{\prime}=1}^{T-1}\sum_{h^{\prime}=g^{\prime}+1}^{T}n_{g^{\prime}}n_{h^{\prime}}(h^{\prime}-g^{\prime})(T-h^{\prime}+g^{\prime})}.$
This weighted-average expression is consistent with the one derived in Goodman-Bacon (2021).
Appendix E Additional Empirical Results
E.1 Impact of Covariates on the Weighted-Average Interpretation
Figure E.1 replicates the analysis from Figure 1 with the inclusion of covariates. Specifically, I include region dummies, log per capita income, and log teen population as covariates. The latter two variables are time-varying, necessitating an adjustment to the weighted-average interpretation. Similar to Figure 1, the FD coefficients $\widehat{\beta}_{\text{FD},k}^{D}$ for $k=1,\ldots,40$ are presented as dots with standard error bars. The TWFE coefficient $\widehat{\beta}_{\text{FE}}^{D}$ is indicated by a dotted horizontal line, while the weighted average of the FD coefficients defined in Theorem 3 is represented by a solid horizontal line. The minimal difference between these two lines suggests that the inclusion of time-varying covariates does not substantially alter the weighted-average interpretation of the TWFE coefficient. In panel (a) of Figure E.1, presenting the analysis for the log teen employment, the TWFE coefficient $\widehat{\beta}_{\text{FE}}^{D}=-0.282$, while the weighted average of the FD coefficients is 0.286-0.286. Panel (b) presents the corresponding analysis for the net job creation rate. The TWFE coefficient $\widehat{\beta}_{\text{FE}}^{D}=-2.80$ remains close to the weighted average of the FD coefficients, which is 2.90-2.90.
E.2 Weights on Causal Effect Parameters
Section 3 provides causal interpretation of the TWFE coefficient, demonstrating that the population TWFE coefficient can be understood as a weighted average of causal effects. For this interpretation to hold, the common trends assumption must be valid for any $k$-period gap, a condition that appears questionable in the minimum wage setting. Even if the common trends assumption is satisfied, a further challenge arises: if some weights are negative and these weights are systematically related to the causal effects of minimum wages, causal interpretation of the TWFE coefficient may be compromised. Therefore, I examine how the weight function derived in Theorem B3 is apportioned across calendar years, states, and minimum wage levels.
Weights on Calendar Years
Figure E.2 illustrates the pattern of weights assigned to different calendar years in the state-year panel data spanning 1979 to 2019. The weights are heavily concentrated in years when the federal minimum wage remained unchanged for extended periods, particularly in the mid-2000s and late 2010s. This pattern aligns with the expectation that years with greater variation in the minimum wages contribute more to the estimation. Although some years are assigned negative weights, their magnitudes are negligible.
Weights on States
Figure E.3 presents the distribution of weights across states, with the vertical axis showing the weight assigned to each state and the horizontal axis indicating the number of years the state’s minimum wage exceeded the federal minimum wage. The analysis reveals a strong positive correlation ($\rho=0.798$) between the state weights and the number of years a state’s minimum wage was above the federal level. This suggests the TWFE coefficient places greater emphasis on states that have experienced more prolonged deviations from the federal minimum wage.
Notably, 24 states are assigned negative weights, most of which closely tracked the federal minimum wage for the majority of the years analyzed. This raises a potential concern if the treatment effects of minimum wage changes differ systematically across states with positive versus negative weights. A common argument in the minimum wage literature posits that minimum wage increases may have negative employment effects only after exceeding the productivity of the marginal job. This threshold is not necessarily reached immediately, especially in labor markets characterized by monopsony power or frictions. If labor productivity differs across states, it could imply heterogeneity in the minimum wage effects. However, this variation would reflect nonlinearity in the treatment effects rather than true effect heterogeneity across states.
Weights on Minimum Wage Levels
Figure E.4 explores whether any specific minimum wage levels receive negative weights in causal interpretation. Two different treatment variables are considered. Panel (a) simply uses the log of the minimum wage as the treatment. Panel (b) uses the deviation of the log minimum wage from the predicted log 25th percentile wage of prime-age (25–54) workers. This alternative treatment variable is motivated by the potential concern raised earlier—if the minimum wage effects differ systematically across states based on how their minimum wages compare to labor productivity levels, then causal interpretation of the TWFE coefficient may be compromised. The predicted log 25th percentile wage is obtained by regressing their observed raw state–year values on state and year indicators, ensuring that this redefinition of the treatment variable does not alter the TWFE coefficient.
Across both panels, the weights assigned to the different minimum wage levels are predominantly positive. This suggests that even when accounting for potential nonlinearity across minimum wage levels, the TWFE coefficient can be interpreted as a meaningful weighted average of the causal effects.
Figure E.1: The TWFE and FD Coefficients with Covariates
(a) Log Employment Rate of Teens
(b) Net Job Creation Rate (%)
  • Notes: The FD coefficients are illustrated with standard error bars. The standard errors are robust to to heteroskedasticity and correlation across observations on the same state. Covariates are census region dummies, log per capita income, and log teen population.
Figure E.2: Weights on Calendar Years
  • Note: The weights on years $t=1,\ldots,T$ are given by $E\left[\omega_{it}^{*}\right]$ using the equation (B.4) and sum to $1$.
Figure E.3: The Relationship Between Weights and State Policies
  • Note: The weights on states $i=1,\ldots,N$ are given by $\sum_{t=1}^{T}\omega_{it}^{*}$ using the equation (B.4) and have a mean of $1$.
Figure E.4: Weights on Treatment Levels, with Two Definitions of Treatment
(a) Log Minimum Wage
(b) Gap Between Minimum Wage and Prime-age 25th Percentile
  • Note: The weights on treatment levels $d\in(\underline{d},\overline{d})$ are given by $\sum_{t=1}^{T}E\left[\psi_{it}^{*}(d)\right]$ using the equation (B.1) and integrate to $1$.